### v1

# Navigating the Future of Federated Recommendation Systems with Foundation Models

Abstract
In recent years, the integration of federated learning (FL) and recommendation systems (RSs), known as Federated Recommendation Systems (FRSs), has attracted attention for preserving user privacy by keeping private data on client devices. However, FRS faces inherent limitations such as data heterogeneity and scarcity, due to the privacy requirements of FL and the typical data sparsity issues of RSs. Foundation models (FMs), such as diffusion models (DMs) and large language models (LLMs), which are widely recognized for producing high-quality content in the image and NLP domains, focus on understanding and mimicking the underlying distribution of the training data. Unlike discriminative models, which learn the boundaries between categories, FMs aim to learn the entire probability distribution of the input data. Thus, the achievements of FMs inspire the design of FRS and suggest a promising research direction: integrating foundation models to address the above limitations. In this study, we conduct a comprehensive review of FRSs with FMs. Specifically, we: 1) summarise the common approaches of current FRSs and FMs; 2) review the challenges posed by FRSs and FMs; 3) discuss potential future research directions; and 4) introduce some common benchmarks and evaluation metrics in the FRS field. We hope that this survey provides the necessary background and guidance to explore this interesting and emerging topic.
Index Terms:
Federated Recommendation System, Foundation Model, Privacy Preserving, Security,
I Introduction
In the digital age, the exponential growth of information has created a need for systems that navigate, filter, and personalize data for individual users. Recommendation Systems (RS) nowadays become crucial tools for filtering online information and helping users discover products, content, and services that align with their preferences [1]. However, the systems traditionally rely heavily on centralized data collection and processing, posing significant privacy risks and operational bottlenecks. The importance of user privacy has never been greater, particularly with stringent data protection regulations such as the General Data Protection Regulation [2] in Europe, which emphasises the need to store user data on theirs local devices, instead of uploading it to central servers. As a novel approach to address these privacy concerns, Google introduced Federated Learning (FL) [3] as a framework designed to train models across decentralized devices, while keeping data localized. This paradigm shift in data processing leverages the computational capabilities of individual devices for local data analysis. Specifically, FL alternates between local model training at the user end and global parameter aggregation from these models on a central server. The integration of FL with RS becomes essential for safeguarding user privacy in recommendation services, which has given rise to the burgeoning field of Federated Recommendation Systems (FRS) [4]. In this field, typically each client consists of a single user’s device. Therefore, unless specified otherwise in this paper, the terms ’user,’ ’client,’ and ’device’ all refer to an individual user. FRS recently have shown promising results in many areas, such as service providing [5, 6], daily scheduling [7], driving planing [8, 9] and more, significantly impacting different facets of daily life. Similar to FL, FRS is required to keep user data local to protect user privacy. However, in most cases, each client typically contains only the data of one user’s accessed item, which is extremely small compared to the total number of items, creating a serious data sparsity problem. In addition, different users have different behaviours and preferences, which can lead to data heterogeneity. The presence of both issues can lead to sub-optimal models and reduced effectiveness.
Figure 1: The foundation model can integrate information contained in data from various modalities during pre-training. The model can then be adapted for a variety of downstream tasks through adaptations such as prompting or fine-tuning.
More recently, the emergence of a novel paradigm for the construction of artificial intelligence (AI) systems has garnered considerable interest in the wake of the remarkable success of ChatGPT [10] and Stable Diffusion [11] in the tasks of language understanding [12, 13, 14, 15] and image generation [16, 17], which we refer to as Foundation Models (FM) [18]. As shown in Fig. 1, this paradigm is built by using self-supervised optimization of training goals to determine how to update model parameters based on model predictions on the large amount of unlabelled training data. This process is referred to as pre-training. Language models, e.g. BERT [19]and RoBERT [20], are usually trained using the next token prediction goal, which refers to the extent to which the model is able to predict the next token in the sequence. One of the most successful examples of language models is ChatGPT, which is based on GPT-3.5. By training on large amounts of text data, it aligns the capabilities of large language models with human intent [21]. Vision models like ViT [22] are typically trained using either contrast learning or diffusion training targets. For contrast learning [23, 24], images are randomly augmented before evaluating the similarity of the model representations. For diffusion models [25, 26], noise is added to the images and the model is gradually de-noised by the target. There are also multi-modal training targets, some of which separate images and text during training, while others consider both. The foundation model, once trained, is used as a plug-in in combination with adaptations to achieve results for a wide variety of downstream tasks. CLIP [27] and DALL·E [28] are two multimodal AI models developed by OpenAI. CLIP focuses on understanding images through natural language, while DALL·E focuses on generating images based on text descriptions. While both models are trained on large datasets of images and text, CLIP is primarily used for image retrieval and classification, while DALL·E excels at generating new images that match text descriptions [29].
As FM is frequently trained using a substantial quantity of data from multiple sources, they are capable of incorporating a considerable amount of additional knowledge when performing a specific downstream task. This feature enables FM to provide diverse training data for task-specific models in downstream task scenarios, thus effectively alleviating the problem of data scarcity. Given these considerations, the application of FM to FRS is not only a viable approach but also holds significant research potential. Such integration will drive innovation in FRS. Although the application of FM to FRS has the potential to be highly beneficial, it must be acknowledged that this field is still in its nascent stages, with an insufficient understanding of the challenges, viable methods, and directions for development. This paper aims to bridge this knowledge gap through an in-depth analysis of the integration of FM and FRS. The article provides a comprehensive examination of the motivations and challenges associated with integrating these two paradigms, with a particular focus on several representative technologies. Additionally, it outlines future development trends and their applications. By elucidating the intersection of FRS and FM, this study aims to promote further exploration and innovation within this emerging field, thereby facilitating its rapid advancement.
Paper Organization. The reminder is structured to provide a thorough examination of the current research landscape, technical challenges, application cases, and future directions of FRS. The organization is as follows: Sec. II outlines the existing surveys within the field of FRS, and distinctly highlights our novel contributions, setting our survey apart from the current literature. Sec. III introduce foundational knowledge in the field of FRS and FM, including definitions, basic architectures, and current scenario classifications. Sec. IV introduces the potential applications and functions of FM according to the three classical phases of FRS. Sec. V delves into the main technical challenges faced by FRS with FM, including data heterogeneity, model aggregation difficulties, and privacy concerns, as well as current methods and their limitations in addressing these challenges. Sec. VII showcases applications of FRS across various domains, such as e-commerce, lifestyle and healthcare, discussing the outcomes and challenges of these real-world implementations. Sec. VIII summarizes the datasets and evaluation methods commonly used in FRS with FM, explaining how to quantify the recommendation performances. Sec. VI explores future research directions in FRS with FM, including unresolved issues and emerging research areas, providing guidance and inspiration for researchers in the field. Sec. IX concludes the survey with a summary of key findings and perspectives, offering insights into the trends and potential future of FRS.
II Related Work and Contribution
In the field of FRSs, various studies have converged on methodologies, privacy preservation, and challenges, albeit with distinct focal points. Yang et al. [30] discuss the practical implementation and evaluation of federated recommendation systems, with a focus on system architecture and algorithmic efficiency. Alamgir et al. [31] provides a comprehensive overview of federated recommendation systems, highlighting techniques, prevailing challenges, and future directions. Javeed et al. [32] focus on the security and privacy concerns in personalized recommendation systems and propose targeted solutions to these challenges. Finally, Sun et al. [33] conduct a survey that compares existing federated recommendation research, highlighting the strengths and limitations of various approaches. All studies emphasize the critical importance of privacy protection and the challenges posed by data heterogeneity and model aggregation. This body of work enhances our understanding of FRSs and provides a foundation for future research to address their inherent challenges. Although a number of surveys [34, 35, 36, 37, 38, 39] focusing on FL with FM already exist, our paper is the first survey work in our current knowledge that focuses on combining FRS with FM.
Contribution. This survey aims to provide a clear theoretical framework for applying FM to FRS and to elaborate on their principles and methods of application. By analyzing existing technologies, this paper strives to drive innovation in integrating the pre-training capabilities of FM into FRS. The article discusses in detail the technical challenges and practical issues faced when integrating FM, such as privacy protection, data heterogeneity, communication efficiency, and model generalization capabilities. It also identifies current research gaps and future directions, aiming to guide subsequent academic research and technological development. Additionally, through application studies, this paper demonstrates how to apply the integration of FM with FRS in practical scenarios, thereby deepening the integration of theory and practice.
III Background and Preliminary
(a) Horizontal FRS
(b) Vertical FRS
(c) Transfer FRS
Figure 2: Typical data categories of FRS.
III-A Federated Recommendation System
Figure 3: A typical framework for a FRS.
Definition 1 (Federated Recommendation System)
A Federated Recommendation System is a technology that uses distributed algorithms for personalized information filtering. It aims to improve the accuracy of information filtering and the effectiveness of personalized services, while maintaining user privacy.
FRS facilitate collaborative learning by pooling analysis and learning capabilities across multiple clients (e.g. users or locations), without the need for direct exchange of raw user data. Fig. 3 illustrates that this type of system typically involves three processes: client model update, communication, and global aggregation. The client model update allows users to train their models locally on their devices using their own data and then upload the updated intermediate parameters to the server. The server performs global aggregation on the parameters sent by all participants, integrating the unique information from each of them. The aggregated parameters are then distributed to the next round of participants, initiating a new training round. The process of uploading and distributing is collectively referred to as communication. These approaches effectively preserves privacy and security.
(a) Centralized FRS
(b) Semi-decentralized FRS
(c) Decentralized FRS
Figure 4: Typical architecture categories of FRS.
III-A1 Taxonomy of FRS
Figure 5: The diagram illustrates the taxonomy of FRS based on two main criteria: Data Distribution and Communication Architecture.
As illustrated in Fig. 5, FRS can be categorized based on different criteria [40, 41]. According to used data distribution, FRS can be divided into Horizontal FRS, Vertical FRS, and Transfer FRS. Each type addresses different data collaboration and learning scenarios. Specifically:
  • Horizontal FRS deals with situations where the feature spaces are similar across different entities, but the sample spaces differ, as shown in Fig. 2a. For instance, two customers who shop at the same store may purchase some of the same products. The horizontal method aggregates model updates from various sources to train recommendation models without sharing raw user data. This approach enhances the accuracy and efficiency of recommendation systems while safeguarding user privacy. Horizontal FRS is currently the most common type, with many studies [42, 43, 44] based on the assumption of this data distribution.
  • Vertical FRS is applied when the feature spaces differ but the sample spaces are similar. This scenario is common when different entities possess different types of data about the same users. An example would be a bank and an online retailer holding distinct perspectives on the same customers, i.e., the bank has credit history while the retailer has shopping history. Vertical FRS [45, 46, 47, 48] trains recommendation models by securely integrating different data features between parties, utilizing richer user information for more accurate personalized recommendations and ensuring data privacy and security.
  • Transfer FRS utilizes principles of transfer learning [49] to transfer knowledge from one domain to another. This type is suited for scenarios with significant differences in both feature and sample spaces across parties. Transfer learning allows the system to leverage data and knowledge from the source domain to enhance recommendation performance in the target domain, even when the target domain lacks sufficient data for independent model training. Transfer FRS [50, 51] is particularly beneficial for emerging markets or user groups with limited data, drawing insights from related domains.
As shown in Fig. 4, FRS can be categorized into centralized, semi-decentralized and decentralized based on their communication architecture. Each type addresses privacy and scalability in different ways:
  • Centralized FRS operates within a FL framework where a central server orchestrates the learning process. User devices (or clients) locally compute updates based on their data and send these updates to the server. The server aggregates these updates to improve the global model, which is then distributed back to the users. While this approach improves privacy by not requiring the sharing of raw data, it still relies on a central authority to manage the model. It effectively addresses privacy concerns by allowing the model to learn from decentralized data sources without centralizing the data itself. This architecture is typically used to overcome data silo issues and improve the performance of RSs without compromising user privacy and data security. Due to its simplicity and intuitive nature, this architecture has become the dominant framework within the field of FRS. There is a great deal of work [52, 43, 44] based on it currently.
  • Semi-decentralized FRS introduces an intermediate layer between the central server and the users, such as edge servers or devices that can perform additional computational or storage tasks. This setup aims to reduce the communication overhead and latency associated with sending updates to a central server, especially in large-scale applications. A specific example is the Semi-decentralized Federated Ego Graph Learning (SemiDFEGL) framework [53], which improves scalability and reduces communication costs by introducing new device-to-device collaborations. It augments local subgraphs with predicted interacted item nodes to exploit high-order collaborative information between users and items in a privacy-preserving manner, which is particularly useful for recommendations based on collaborative filtering and graph neural networks.
  • Decentralized FRS, which employs peer-to-peer communication architecture [54], distributes the learning process completely to all participating devices, without the need for a central server for model aggregation. This approach maximizes privacy and data ownership, but presents challenges in coordinating model updates and ensuring model convergence. Zheng et al. [55] proposed a FRS model called DGREC by adopting a decentralized graph neural network, which constructs a local intra-item hypergraph and a global inter-user graph for each user, allowing users to freely choose whether to disclose their interactions. Li et al. [56] introduced DFedRec, which uses a privacy-aware client-level structured graph to share model parameters only with relevant neighboring users, thereby reducing communication costs and protecting user privacy.
Each of these architectures offers a trade-off between privacy, communication efficiency and the degree of decentralization. Centralized systems simplify model aggregation, but rely on a central authority. Semi-decentralized systems aim to balance efficiency and privacy with an intermediate layer, while decentralized systems offer the highest level of privacy at the cost of more complex model coordination.
III-B Foundation Model
The rapid increase in the performance of computer hardwares, e.g., GPUs, the increasing maturity of transformer architectures, and the public availability of large amounts of training data have been three key factors in the emergence of FM [18]. According to the work of Stanford HAI [18], we have the following definition of FM:
Definition 2 (Foundation Model)
A foundation model is defined as any model trained on extensive data (typically using large-scale self-supervised learning) that is capable of adapting to a wide range of downstream tasks, for instance, through fine-tuning.
In recent years, the scale and scope of FM have greatly expanded our imagination of potential applications. Such models typically have billions or even trillions of parameters, allowing them to learn more complex patterns and knowledge. They can be adapted to new tasks through fine-tuning or zero-sample learning. For instance, the GPT-3 model [57], comprising 175 billion parameters, is capable of adapting and perform various tasks with the aid of natural language prompts, despite the fact that a significant proportion of these tasks have not been explicitly trained. FM is distinguished by two key characteristics: 1) Emergence, which refers to the implicit induction of system behavior from examples, as opposed to explicit design; and 2) Homogenization, which indicates that the method of constructing machine learning systems tends to be unified across a wide range of applications. Although FM is based on standard deep learning and transfer learning, their scale brings new emergent abilities, and their effectiveness in many tasks has motivated homogenization. While homogenization provides strong leverage, it also requires caution, as defects in foundation models can be inherited by all downstream adapted models. Moreover, despite the broad deployment prospects of foundation models, due to their emergent properties, our understanding of how they work, when they fail, and what they can actually do is still quite limited.
Figure 6: The diagram illustrates the taxonomy of FM based on two main criteria: Data Type and Functionality.
III-B1 Taxonomy of FM
As shown in Fig. 6, similar to FRS, FM can also be classified based on data type or functionality. Specifically, based on the type of data used during training, FM is mainly divided into followings [18]:
  • Language FM: These types of FM mainly deal with textual data and are trained on large textual datasets to understand and generate language. They are very effective in natural language processing (NLP) tasks such as machine translation, text summarization, and sentiment analysis. Typical language models include BERT [19], GPT-3 [57], and T5 [58], which typically use deep transformer architectures that are able to capture complex linguistic regularities and demonstrate excellent performance on multilingual tasks.
  • Vision FM: Vision FM focuses on processing and understanding image data. These models are able to perform tasks such as object recognition, image segmentation, and visual reasoning by learning large amounts of image data. For example, DINOv2 [59] and SAM [60] are basic models trained specifically for visual tasks, and using self-supervised learning methods, these models learn valid visual representations without labelling the data.
  • Multimodal FM: Multimodal models can simultaneously process and understand multiple types of data, such as text and images. Such models can excel in cross-modal tasks such as image captioning and visual quizzing by integrating information from different modalities. CLIP [27] and DALL·E [28] are representative of such models that are able to understand the relationship between an image and its corresponding textual descriptions, demonstrating flexibility and robustness in handling multiple data types.
Functionally, FMs are generally categorized into two types [18, 61]:
  • Discriminative FM: The main task of discriminative FM is to distinguish or predict specific outputs from given input data. These models are generally based on the BERT family and are more concerned with learning decision boundaries from the data to perform classification, regression, or other predictive tasks.
  • Generative FM: The core goal of generative FM is to learn the distribution of the data and be able to generate new data samples. These models, such as GPT-3, DALL·E, etc., are able to capture the underlying structure of the data and thus generate new instances that are similar but different from the training data. Generative FM has a wide range of applications in areas such as text generation and image generation.
III-B2 Adapter
Fig. 1 illustrates a commonly used technique for pre-trained models, known as the Adapter. This technique adds new lightweight layers while maintaining the original parameters of the pre-trained model, enabling fine-tuning and extension for specific tasks. This approach is suitable for multi-task learning and tasks performed in resource-limited environments. Adapters, such as LoRA [62] and QLoRA [63], have found wide applications in natural language processing [62] and computer vision tasks.
IV Federated Recommendation Systems with Foundation Models
As described earlier in Fig. 3, a typical FRS typically consists of three stages: client model update, communication, and global aggregation. Integration with FM should also occur at these three stages.
IV-A Client Model Update
In FRS, the clients have the following characteristics:
  • As previously described, the data needed for model training in FRS, such as user information and interaction history, are privacy-sensitive and are therefore required to stay on the user’s device, creating data silos.
  • Moreover, each user’s data is minimal compared to the total dataset, and each user only accesses a small portion of the item set, leading to data sparsity.
  • The data on the client is influenced by user preferences, such as user behavior and product preferences. The resulting data distribution often does not meet the independent and identically distributed (IID) assumption, leading to the challenge of data heterogeneity.
  • Additionally, user devices are generally consumer products like mobile phones and personal computers, characterized by unstable communication and limited computational resources. This requires that the computational load of the models deployed on the client and the amount of information exchanged with the server be kept as low as possible.
FM is typically pre-trained on large, diverse datasets, acquiring a broad range of features and knowledge. This pre-training endows them with prior knowledge that allows for rapid adaptation to specific client data through fine-tuning. Consequently, these models can learn general feature representations, providing a certain level of adaptability to different data distributions. Therefore, clients can effectively fine-tune the model with a minimal amount of local data, achieving good performance on specific downstream tasks. Additionally, by fine-tuning foundation models locally, sensitive data does not need to leave the device, thereby enhancing data privacy.
However, it is important to note that when FM is applied to data significantly different from the training distribution, performance degradation may occur. This issue, known as out-of-distribution generalization, represents a challenging aspect that FMs need to overcome. Moreover, if biases exist in the training dataset, FMs might learn and amplify these biases, leading to unfair outcomes across different data distributions. Although fine-tuning FMs requires significantly fewer resources than training from scratch, it usually still demands substantial computational resources for effective fine-tuning and updating. This requirement could limit their application on resource-constrained clients, particularly in FRS, where each user represents a client, thus potentially restricting the deployment of FM due to limited computational resources.
IV-B Communication
Communicate Efficiency. Similar to FL, FFRS involves significant data transfer between numerous clients and a central server, making communication efficiency a critical factor in the duration of the entire learning process. To enhance communication efficiency, model compression techniques such as quantization or sparsification of model parameters can be utilized to reduce the volume of data transmitted. Periodic averaging is another strategy that reduces communication overhead by decreasing the frequency of model parameter uploads.
FM, due to its parameter sharing and hierarchical features, allow clients to transmit only fine-tuned parameter updates rather than the entire model’s parameters. This approach can significantly reduce communication overhead while maintaining or even enhancing the model’s performance on specific tasks. Similarly, since foundation models have already learned rich feature representations during the pre-training phase, they can adapt more quickly to new tasks. This means that in FRS, clients may achieve satisfactory performance with fewer iterations, thus reducing the frequency of communication.
Client Selection. In each iteration of FRS, considering resource limitations and privacy concerns, the central server may choose to communicate only with a subset of clients. Research indicates that clients should not participate in consecutive training rounds as attackers could potentially infer information about the client from transmitted gradients or model parameters. Client selection strategy determines which clients will participate in the current round based on various factors such as computational power, network stability, data diversity, and quality. Client selection can also be conducted through random sampling or based on statistical characteristics of the data to ensure the model learns from diverse data sources, enhancing its generalization ability.
Since FM already possesses a broad knowledge base and generalization capabilities, when selecting clients for model training, greater emphasis can be placed on data compatibility and coverage. Specifically, priority can be given to clients that can provide data types less represented or missing in the pre-training phase, thereby supplementing and enhancing the model’s performance in these areas. For instance, clients whose data can significantly improve the model’s performance on specific tasks can be considered more valuable. Incentive mechanisms can be set up to encourage the participation of these high-value data providers in the training process. However, it is important to ensure that the same client does not participate in consecutive training rounds to minimize the risk of privacy breaches.
Privacy Protection. A major advantage of federated learning is the ability to train models without compromising user privacy, making privacy protection a primary consideration in designing communication protocols. Privacy risks are mitigated by transmitting model updates, such as gradients or parameters, instead of raw data from clients to the server. Additionally, technologies like homomorphic encryption, secure multi-party computation (SMC), or differential privacy can further enhance privacy protection during communication.
IV-C Global Aggregation
Global aggregation in FRS integrates model parameters independently trained by various clients to form a unified global model. This process not only enhances the accuracy and generalization of the recommendation system but also strengthens the system’s robustness by protecting client data privacy, optimizing resource usage, ensuring model synchronization and fairness, thereby improving recommendation performance and efficient resource utilization without compromising user privacy.
IV-C1 Typical Average
The Typical Average scheme, proposed with FedAvg [3], is the most common aggregation strategy where the server computes a simple arithmetic average of the model updates received from client devices. It’s popular due to its simplicity and effectiveness in many scenarios
IV-C2 Weighted Average
In the Weighted Average strategy, model updates are weighted based on certain criteria such as the volume of data on each client or the reliability of the data source, allowing more significant contributions to have a proportionally greater impact on the global model. This approach can be more effective than the typical average scheme, especially in non-IID data environments, where data distribution varies significantly across devices
IV-C3 New Future Methods
When integrating FM, exploring new aggregation technologies becomes particularly important to address the challenges posed by the scale and complexity of the models. Beyond the traditional weighted average method, some emerging aggregation strategies have shown potential in managing these challenges. For instance, the technique called Model Soups enhances model accuracy and robustness by averaging the weights of models that have been fine-tuned with different hyperparameters. Additionally, aggregation strategies based on the Mixture of Experts (MoE) utilize multiple specialized sub-models, each optimized for specific tasks or data types. These strategies dynamically adjust model weights based on data-driven methods, allowing for flexible adjustment of the aggregation process in response to real-time data changes, thus improving the overall model performance and adaptability. These emerging technologies not only improve the effectiveness of aggregation but also better address the challenges from large-scale distributed data, making them valuable aggregation strategies worth further research and application in the field.
Additionally, leveraging machine learning techniques to dynamically adjust aggregation methods based on real-time data quality and integrity assessments could improve both the robustness and efficiency of FRS.
V Challenges
The integration of generative models into federated recommendation systems heralds a new frontier in personalized content delivery, promising to enhance user experiences with tailored suggestions while respecting privacy concerns. However, this promising union is not without its complexities. As we venture further into this domain, a myriad of challenges emerges, each posing unique hurdles that must be carefully navigated. These challenges span the spectrum from data heterogeneity and privacy concerns to communication overheads and resource-intensive computations. Understanding and addressing these issues is paramount to the successful deployment and operation of federated recommendation systems that are both effective and secure. In this section, we delve into each of these challenges, exploring their implications and discussing potential strategies for overcoming them.
V-A Data heterogeneity
The diversity in user-generated data can complicate the generative model’s ability to identify a common representation that accurately reflects all users’ preferences. This challenge necessitates the development of robust mechanisms capable of handling varied data inputs and providing nuanced, precise recommendations. Techniques such as transfer learning and multi-task learning can be employed to better adapt to the heterogeneity of data.
V-B Privacy and security risks
The collection of user data is essential for refining recommendation algorithms, but it also poses significant privacy and security risks. While federated learning offers a solution by processing data locally, generative models must be carefully designed to prevent the leakage of sensitive information through their generated outputs. Privacy-preserving methods like differential privacy and secure multi-party computation can be integrated to enhance data security.
V-B1 Memeber Inference Attacks
Membership inference attacks aim to determine whether specific data samples were used in training a model. Such attacks can reveal the presence of sensitive data, for example, by disclosing whether a patient’s medical records were used to train a disease prediction model. Under the FL setting, attackers might use model updates obtained from various participants to infer which data were used for training. Particularly in FRS with FM settings, the complexity and depth of the models make membership inference attacks more covert, thereby complicating defense efforts.
V-B2 Data Reconstruction Attacks
Data reconstruction attacks aim to rebuild or approximate the actual data used in training. This is typically achieved through optimization techniques such as model inversion and gradient matching, where attackers attempt to generate data samples similar to the training data by accessing the model’s outputs or gradients. Under the FL setting, although the original data does not leave the owner’s device, aggregated updates may still leak sufficient information to enable attackers to reconstruct the original data. Particularly when large-scale foundation models are used, this risk may be exacerbated due to the models’ high capacity for data representation.
V-B3 Poisoning Attacks
In FL, poisoning attacks represent a security threat aimed at disrupting or manipulating the learning process and outcomes through malicious modifications to data or model parameters. These attacks primarily take two forms: data poisoning attacks and model poisoning attacks. Specifically:
  • Untargeted Attacks: Untargeted attacks aim to disrupt the entire training process, causing the global model to fail to converge or significantly degrade in performance. These attacks are typically executed by injecting noisy data or incorrect information into the model training process. In a FL environment, attackers might submit model updates containing erroneous gradients or parameter updates, thereby disturbing the learning process of the aggregated global model. The challenge of these attacks lies in the ability of the attackers to pass the poisoned data through the FL system’s normal aggregation process without detection.
  • Targeted Attacks: In contrast, targeted attacks aim to cause the model to produce incorrect outputs for specific inputs, without affecting the model’s performance on most other inputs. This type of attack is usually carried out by injecting a small amount of carefully designed poisoned samples into the training data. These samples contain specific triggers that, when encountered by the model, lead to predetermined incorrect outputs. This attack method is particularly covert in FL, as attackers can embed these triggers during local training, and these modifications may only affect a small part of the model, making them difficult to detect in the global model.
V-C Communication Cost
Federated systems necessitate ongoing communication between users and servers to synchronize model updates. The large parameter size of generative models can significantly increase the demand for communication resources. To address this, efficient communication protocols and model compression techniques can be utilized to reduce the bandwidth required for updates.
V-D Data Scarcity
The scarcity of user data is a significant challenge, particularly for new or less active users, as it can hinder the model’s ability to learn accurate preferences. To combat this, data augmentation strategies and synthetic data generation can be employed to enrich the training data and improve model generalization.
V-E High Resource Demand
Generative models are computationally demanding, requiring substantial storage and processing capabilities. This can be particularly challenging in resource-constrained environments, such as mobile devices. Optimizing model architectures and leveraging distributed computing can help alleviate the strain on resources.
V-F Problematic Synthetic Data Generated
The quality of synthetic data generated by generative models is crucial. Low-quality synthetic data can negatively impact the model’s performance if used in training. Ensuring the synthetic data closely resembles real data distributions is essential for maintaining the integrity of the recommendation system.
V-G Robustness Consideration
Generative models within federated recommendation systems must be robust against data anomalies and capable of maintaining stable performance even when faced with shifts in data distribution. This requires the implementation of robust learning algorithms that can adapt to changes in the data landscape without significant degradation in performance.
VI Future Directions
This section discusses applications that highlight the great potential of generative models in federated recommender systems. These applications not only have the potential to improve recommendation quality, but also to provide personalised services while protecting user privacy. As technology continues to advance, we can expect generative models to play an increasingly important role in future FRSs.
VI-A Personalized Content Generation
Generative models such as VAEs and GANs can be used to generate personalized user profiles within federated recommendation systems. These models can capture latent features of user preferences and generate recommendations that reflect these characteristics, providing more accurate personalized suggestions.
VI-B Privacy Protection
Privacy is paramount in federated recommendation systems. Generative models can protect user data privacy by generating representations of user preferences locally and then only sending these representations to a central server for recommendations, without disclosing raw user data.
VI-C Data Augmentation
Diffusion models and large language models can be used to generate additional training data, which is particularly useful in federated learning environments where data sparsity is often an issue. By generating synthetic data, the diversity of the training set can be increased, enhancing the performance of the recommendation system.
VI-D Cross-Domain Recommendations
Federated recommendation systems allow multiple organizations to share in the training of a recommendation model without sharing user data. Generative models can be used within this framework to generate cross-domain recommendations, sharing generated latent user features across different organizations rather than raw data.
VI-E Cold Start Problem Solution
For new users or new items, traditional recommendation systems face the so-called ”cold start” problem. Generative models can predict the latent features of these new entities and generate initial recommendations, quickly initiating the recommendation process.
VI-F Enhanced Recommendation Explainability
Large language models can be used to generate explanations for recommendations, helping users understand why a particular piece of content is recommended. This is crucial for increasing user trust and acceptance of recommendation systems.
VI-G Real-Time Recommendations
Generative models can quickly produce recommendations, which is useful for recommendation scenarios that require real-time feedback, such as news recommendations or real-time event recommendations.
VI-H Reduction of Recommendation Bias
By generating a diverse range of recommendations, generative models can help reduce biases in recommendation systems, providing more equitable recommendation outcomes.
VI-I System Robustness Enhancement
In federated recommendation systems, generative models can be used to simulate various user behaviors and preferences, thereby testing and improving the robustness of the recommendation system.
VI-J Multimodal Recommendations
Combining multiple types of data such as images, text, and audio, generative models can be used to create multimodal recommendations, such as recommending travel destinations that combine images with textual descriptions.
VII Applications
VII-A eHealthcare Systems
eHealthcare systems leverage modern technology to enhance the quality and efficiency of medical services, such as remote diagnosis and patient monitoring. Federated recommendation systems allow different medical institutions to share insights from patient data without sharing the data itself, which is crucial for protecting sensitive health information. Generative models use this data to provide doctors and patients with personalized treatment plans and health advice, improving diagnostic accuracy and treatment outcomes. Thus, federated recommendation systems combined with generative models can integrate patient health data across institutions to create complex treatment pathways and simulate disease progression, enabling doctors to better predict patients’ responses to specific treatments and provide more precise medical services.
VII-B Industrial Applications
In manufacturing, automation and intelligent technologies significantly enhance production efficiency and safety. Federated recommendation systems foster technological collaboration between different enterprises by sharing non-sensitive data, accelerating product design and optimization processes. This not only shortens product development cycles but also protects corporate trade secrets. Generative models can design new mechanical parts or product prototypes by simulating different design variables, quickly proposing multiple design solutions; they can also help predict machine failures, perform maintenance in advance, and reduce downtime, thereby enhancing production efficiency. Federated recommendation systems combined with generative models enable the sharing of improvements and innovations while maintaining data privacy between companies.
VII-C Consumer Services
Analyzing consumer behavior helps businesses deeply understand customer needs. Federated recommendation systems can integrate data from different service providers while protecting user privacy, achieving personalized services. Generative models provide personalized recommendations based on this data, enhancing user experience. For example, in travel recommendation systems, models generate customized travel itineraries based on users’ historical preferences. By simulating different consumer behavior patterns, businesses can predict market trends, adjust service strategies, and enhance competitiveness. Federated recommendation systems combined with generative models analyze user behavior data from different service providers to generate service plans that meet individual needs, promoting personalized service development.
VII-D Retail and E-commerce
Applying recommendation systems to online shopping and retail provides a personalized shopping experience. Federated recommendation systems in the retail and e-commerce sectors analyze consumer shopping habits to generate personalized product recommendations, increasing users’ purchasing desire and satisfaction. Retailers and e-commerce platforms can use generative models to predict consumer buying behavior, thereby offering more accurate product recommendations, optimizing inventory, and enhancing sales efficiency. Federated recommendation systems combined with generative models allow retailers and e-commerce platforms to consider market conditions comprehensively, predicting consumer buying behavior and providing more precise product recommendations, optimizing inventory, and enhancing sales efficiency.
VII-E Finance and Banking
In the financial sector, recommendation systems use data analysis and prediction models to optimize financial services, applicable for credit assessments, stock market analysis, etc., offering clients tailored financial products. Federated recommendation systems ensure transaction data privacy while offering clients tailored financial products. Generative models can simulate market behaviors, predicting stock prices or financial product returns. These models help financial institutions formulate more accurate investment strategies while reducing risks. In finance, federated recommendation systems combined with generative models help financial institutions protect client privacy while simulating market changes and user behavior, offering advice on investment products and credit strategies, thereby enhancing clients’ investment returns and satisfaction.
VII-F Smart Cities
In smart city applications, systems need to process and analyze large-scale urban data. From optimizing traffic flow to devising public safety strategies, federated recommendation systems can help city planners use resident data without directly sharing it to recommend strategies for improving public services and infrastructure, enhancing urban management efficiency and improving residents’ quality of life. Additionally, using generative models to simulate the effects of different strategies can find optimal solutions for managing and allocating urban resources. For example, by simulating traffic flow, models can predict congestion and propose solutions, improving urban traffic efficiency. Federated recommendation systems combined with generative models can help city planners optimize resource allocation and management efficiency without sharing resident data.
VII-G Social Networking
Social platforms facilitate interpersonal communication and information sharing online, using recommendation systems to suggest content to users, such as friends or topics of interest, to increase user engagement. Federated recommendation systems can enhance user engagement and satisfaction while protecting user privacy. Generative models can generate personalized content recommendations, enhancing user engagement and satisfaction while ensuring the protection of user privacy, increasing platform attractiveness and user retention.
VII-H Education
In the field of education, using recommendation systems can improve teaching methods by making appropriate recommendations based on students’ progress and abilities, providing a personalized learning experience. Generative models can generate personalized learning materials and courses based on students’ progress and interests. These customized learning resources can enhance students’ learning efficiency and interest. Federated recommendation systems combined with generative models can generate personalized learning paths based on students’ feedback and learning history. By simulating teaching scenarios and student learning processes, these systems can provide resources that better meet individual learning needs.
VIII Resources
VIII-A Datasets
Figure 7: Two types of Feedbacks.
TABLE I: Datasets with Explicit Feedback.
Context
Dataset #User #Item #Feedback Sparsity User Item Feedback Timestamp
Amazon Beauty 1,210,271 249,274 2,023,070 99.99%
Amazon Books 8,026,324 2,330,066 22,507,155 99.99%
Amazon CDs 1,578,597 486,360 3,749,004 99.99%
Amazon Cell Phone 2,261,045 319,678 3,447,249 99.99%
Amazon Clothing 3,117,268 1,136,004 5,748,920 99.99%
Amazon Games 826,767 50,210 1,324,753 99.99%
Amazon Garden 714,791 105,984 993,490 99.99%
Amazon Home 2,511,610 410,243 4,253,926 99.99% $\checkmark$ $\checkmark$
Amazon Kindle 1,406,890 430,530 3,205,467 99.99% $\checkmark$ $\checkmark$
Amazon Movies 2,088,620 200,941 4,607,047 99.99% $\checkmark$ $\checkmark$
Amazon Music 478,235 266,414 836,006 99.99% $\checkmark$ $\checkmark$
Amazon Sports 1,990,521 478,898 3,268,695 99.99% $\checkmark$ $\checkmark$
Amazon Toys 1,342,911 327,698 2,252,771 99.99% $\checkmark$ $\checkmark$
Amazon Reviews Amazon Video 426,922 23,965 583,933 99.99% $\checkmark$ $\checkmark$
ML-100K 943 1,682 100,000 93.70% $\checkmark$ $\checkmark$ $\checkmark$
ML-1M 6,040 3,952 10,000,209 95.81% $\checkmark$ $\checkmark$ $\checkmark$
ML-10M 69,878 10,681 10,000,054 98.69% $\checkmark$ $\checkmark$
MovieLens ML-20M 138,493 27,278 20,000,263 99.47% $\checkmark$ $\checkmark$
Yelp2018 1,326,101 174,567 5,261,669 99.99% $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
Yelp2020 1,968,703 209,393 8,021,122 99.99% $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
Yelp2021 2,189,457 160,585 8,635,403 99.99% $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
Yelp2022 1,987,929 150,346 6,990,280 99.99% $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
Yelp Yelp-full 5,556,436 539,254 28,908,240 99.99% $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
Anime 73,515 11,200 7,813,737 99.05%
Book Crossing 105,284 340,557 1,149,780 99.99% $\checkmark$ $\checkmark$
Douban 738,701 28 2,125,056 89.73% $\checkmark$ $\checkmark$
Epinions 116,260 41,269 188,478 99.99% $\checkmark$ $\checkmark$
Goodreads 876,145 2,360,650 228,648,342 99.99% $\checkmark$ $\checkmark$
Jester 73,421 101 4,136,360 44.22%
Netflix 480,189 17,770 100,480,507 98.82% $\checkmark$
Yahoo Music 1,948,882 98,211 11,557,943 99.99% $\checkmark$
TABLE II: Datasets with Implicit Feedback.
Context
Dataset #User #Item #Feedback Sparsity User Item Feedback Timestamp
MIND_large_train 711,223 27,047 83,507,374 99.57% $\checkmark$
MIND_large_dev 255,991 6,998 14,085,557 99.21% $\checkmark$
MIND_small_train 50,001 20,289 5,843,444 99.42% $\checkmark$
MIND MIND_small_dev 50,001 5,370 2,740,998 98.98% $\checkmark$
QK-video 5,022,750 3,753,436 142,321,193 99.99% $\checkmark$ $\checkmark$
QK-article 1,325,838 220,122 46,111,728 99.98% $\checkmark$ $\checkmark$
QB-video 34,240 130,647 1,701,171 99.96% $\checkmark$ $\checkmark$
Tenrec QB-article 24,516 7,355 348,736 99.81% $\checkmark$ $\checkmark$
Adressa 15,514 923 2,717,915 81.02% $\checkmark$ $\checkmark$
Foursquare 1,083 38,333 227,428 99.45% $\checkmark$ $\checkmark$
Gowalla 196,591 950,327 6,442,890 99.99%
Last.FM 1,892 17,632 92,834 99.72% $\checkmark$
Pinterest 55,187 9,911 1,445,622 99.74%
Steam 2,567,538 32,135 7,793,069 99.99% $\checkmark$ $\checkmark$ $\checkmark$
TaFeng 32,266 23,812 817,741 99.89% $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
Tmall 963,923 2,353,207 44,528,127 99.99% $\checkmark$ $\checkmark$
In FRS, the datasets used can be categorized into two types based on user feedback: explicit feedback and implicit feedback, as shown in Fig. 7. Explicit feedback includes direct responses from users about their preferences. This typically includes ratings (such as a 1-5 scale), user comments, and like/dislike statements. Such data is clear and provides straightforward insights into user preferences, making it highly valuable for training recommendation models. Common datasets with explicit feedback include the following shown in Table ?????????I:
  • Amazon Reviews [^1] https://cseweb.ucsd.edu/ jmcauley/datasets.html#amazon_reviews: The Amazon Reviews dataset is a large-scale dataset that contains product information across various categories such as Books, CDs, and Music. It includes reviews (ratings, text, helpfulness votes) and product metadata (description, category information, price, brand, and image features). There are three updated versions of this dataset from the years 2014 [64, 65], 2018 [66], and 2023 [67].
  • MovieLens Datasets [^2] https://grouplens.org/datasets/movielens/: The MovieLens datasets [68], initially released in 1998, capture individuals’ stated movie preferences. These preferences are recorded as tuples, with each tuple showing a person’s rating (from 0 to 5 stars) for a movie at a specific time. Users enter these ratings through the MovieLens website, which provides personalized movie suggestions based on these ratings.
  • Yelp Datasets [^3] https://www.yelp.com/dataset: This dataset is a subset of Yelp’s business, review, and user data. It was initially developed for the Yelp Dataset Challenge, which allows students to study or analyze Yelp data and present their insights. In total, there are four versions of the Yelp datasets.
  • Anime [^4] https://www.kaggle.com/datasets/CooperUnion/anime-recommendations-database: This dataset collects user preference data from the MyAnimeList website. It contains information from 73,516 users on 12,294 anime titles. Users can add anime to their completed list and rate them, and this dataset compiles these ratings.
  • Book Crossing [^5] https://grouplens.org/datasets/book-crossing/: The Book-Crossing dataset is a well-structured collection of data collected by Cai-Nicolas Ziegler in a 4-week crawl from the Book-Crossing community. This dataset primarily comprises user interactions that include book ratings, ranging from 0 to 10.
  • Douban [^6] https://www.kaggle.com/datasets/utmhikari/doubanmovieshortcomments: The Douban Movie dataset is a Chinese website where internet users can post their opinions and comments about films. This dataset contains over 2 million short comments on 28 movies from the Douban Movie website.
  • Epinions [^7] https://cseweb.ucsd.edu/ jmcauley/datasets.html#social_data [69]: This dataset was collected from Epinions.com, a popular online consumer review site. It includes trust relationships between users and covers a period from January 2001 to November 2013.
  • Goodreads [^8] https://www.kaggle.com/datasets/jealousleopard/goodreadsbooks: This dataset includes reviews from the book review website Goodreads, along with various attributes describing the books. Importantly, the dataset captures different levels of user interaction, from adding books to a shelf, to rating them, to reading them.
  • Jester [^9] https://eigentaste.berkeley.edu/dataset/: The Jester dataset focuses exclusively on jokes. Users of the Jester online platform rate jokes and these ratings are then used to personalize joke recommendations for them.
  • Netflix [^10] https://www.kaggle.com/datasets/netflix-inc/netflix-prize-data: Netflix provided a training dataset consisting of 100,480,507 ratings from 480,189 users for 17,770 films. Each rating is represented as a set of four elements: ¡user, movie, rating date, rating score¿. The user and the movie are identified by integer IDs, and the rating scores range from 1 to 5 stars, also as integers.
  • Yahoo Music [^11] https://webscope.sandbox.yahoo.com/catalog.php?datatype=r: The Yahoo Music dataset is known for its large scale and diversity. It contains a large collection of user ratings on different musical elements such as tracks, albums, artists and genres. This dataset was used in the KDD-Cup 2011 competition, where participants were asked to analyse user preferences in music based on these ratings.
Implicit feedback, on the other hand, is derived from user actions that indirectly indicate preferences, such as bookmarks, video/music play history, or click-throughs. Although implicit feedback does not directly express user likes or dislikes, it is rich and captures user behaviour more comprehensively.
  • MIND [^12] https://msnews.github.io/ [70]: The MIND dataset, sourced from the Microsoft News website, is a large-scale collection of approximately 160,000 English news articles and over 15 million user interaction records. It has been designed to advance research in news recommendation systems. It includes detailed textual content for each story and anonymized user interaction data to ensure privacy.
  • Tenrec [^13] https://github.com/yuangh-x/2022-NIPS-Tenrec [71]: The Tenrec dataset is a comprehensive benchmark dataset for RSs, featuring user interactions from two recommendation platforms across four dataset files: QK-video and QB-video for video actions, and QK-article and QB-article for article actions.
  • Adressa [^14] https://reclab.idi.ntnu.no/dataset/ [72]: The Adressa dataset is a corpus of Norwegian news articles related to anonymous users. It is a collaborative project between the Norwegian University of Science and Technology and Adressavisen. The objective is to gain insight into the nature of news articles and their readers.
  • Foursquare [^15] https://sites.google.com/site/yangdingqi/home/foursquare-dataset [73]: This dataset comprises check-in data from New York City, collected over a period of approximately ten months (from 12 April 2012 to 16 February 2013). It encompasses 227,428 check-ins in New York City, with each check-in recorded with its timestamp, GPS coordinates, and detailed venue category.
  • Gowalla [^16] https://snap.stanford.edu/data/loc-gowalla.html [74]: Gowalla is a location-based social networking website where users can post their whereabouts by checking in. The dataset comprises data collected from the public API, which represents an undirected friendship network with 196,591 nodes and 950,327 connections. Additionally, it records 6,442,890 check-ins made by these users between February 2009 and October 2010.
  • Last.FM [^17] https://grouplens.org/datasets/hetrec-2011/ [75]: The Last.FM dataset represents a valuable resource that has been extensively utilized in the field of music information retrieval and RSs. It captures detailed information regarding music listening events from users. Each listening event is further enhanced with user demographics and specific descriptors that reflect their music tastes and consumption behaviours.
  • Pinterest [^18] https://github.com/edervishaj/pinterest-recsys-dataset [76]: The Pinterest dataset represents a valuable resource for a variety of research and analytical purposes. It encompasses a diverse range of data, including images, user features, interests, and user interactions.
  • Steam [^19] https://github.com/kang205/SASRec [77]: The Steam dataset is a collection of information about games published on the Steam platform. It includes details such as game names, release dates, genres, developers, publishers, and other relevant information.
  • TaFeng [^20] https://www.kaggle.com/datasets/chiranjivdas09/ta-feng-grocery-dataset: The TaFeng dataset is a comprehensive collection of supermarket shopping data, including detailed transaction records from the Ta Feng supermarket in Taiwan, covering a period from November 2000 to February 2001. The dataset comprises a variety of data points, including customer demographics, product categories, and detailed item descriptions along with quantities purchased.
  • Tmall [^21] https://tianchi.aliyun.com/dataset/53[78]: The Tmall dataset is a comprehensive collection from Tmall, comprising anonymized user shopping records over a six-month period up to and including the ”Double 11” event. It should be noted that the data is selectively sampled to address privacy concerns.
Both types of feedback play a critical role in the development of FRSs, providing diverse insights into user preferences that help improve the accuracy and relevance of the recommendations provided.
VIII-B Metrics
In the field of FRSs, the use of evaluation metrics is fundamental to assessing and refining the performance of our algorithms. Metrics serve as a quantitative lens through which we can observe how closely the system’s suggestions match users’ actual interests and preferences.
For predicting how well a system can estimate user preferences, there are several metrices measuring the prediction errors to judge the accuracy of the predictions, such as Mean Absolute Error (MAE), Mean Squared Error (MSE) and Root Mean Squared Error (RMSE). When it comes to classifying items, i.e., determining whether a user will like a product or not, we look at metrics such as Precision, Recall, Hit Ratio (HR), F1 Score, Accuracy and AUC. These tell us how correctly the RS is classifying items, and the F1 score helps us balance the Precision and Recall.
Then to measure the item ranking ability, which is about listing recommendations in the right order, Average Precision (AP) and Mean Average Precision (MAP) are key to this, as they assess the quality of the order of recommendations. Metrics such as Mean Reciprocal Rank (MRR), Normalized Mutual Rank (NMR) and Normalised Discounted Cumulative Gain (NDCG) also contribute by assessing how well the top recommended items are ranked.
On a broader scale, we consider recommendation-centric metrics such as Diversity, which ensures that a variety of items are suggested, and Coverage, which measures how many items from the catalogue are recommended. There are also some user-centric metrics include Novelty, which measures how new or surprising the recommendations are, and Degree of Agreement (DOA), which quantifies the level of concordance between the ranking of items produced by a recommendation system and the ranking preferred by the user. Moreover, business metrics such as Click-Through Rate (CTR) are critical to assessing the system’s impact on user engagement and the company’s bottom line, and Conversion Rate (CVR) measures how efficient an algorithm is at providing recommendations that lead to user purchases. There also are some metrics for measuring other functionality. For example, Gini Index evaluates the fairness of recommendation distribution, with lower values indicating more equitable distribution across items.
Furthermore, in the generative recommendation scenario, FMs such as DMs and LLMs can generate items that have never appeared in historical data and recommend them to users. In this case, how to evaluate the generative recommendation capability of these generative FMs remains an open question.
IX Conclusion
This study comprehensively examines the integration of FRSs with FMs, a direction that has gained attention for its ability to protect user privacy. The article begins by summarizing common approaches of FRSs and FMs, then delves into the challenges faced during integration. To address these challenges, the paper proposes various strategies including using transfer learning and multitask learning techniques to adapt to data diversity, employing privacy-preserving methods like differential privacy and secure multi-party computation, and reducing communication overhead through model compression and efficient communication protocols.
This work also discusses the future research directions directions indicating that FRSs can provide more accurate personalized recommendations while better protecting user privacy. Additionally, the paper showcases applications of FRS in various fields, demonstrating their potential and value in the real world. Through this study, we aim to provide theoretical guidance for integrating FRS with FM, directing future research and technological advancements to collectively advance this field.