### v2

# Navigating the Future of Federated Recommendation Systems with Foundation Models

Abstract
In recent years, the integration of federated learning (FL) and recommendation systems (RS), known as Federated Recommendation Systems (FRS), has attracted attention for preserving user privacy by keeping private data on client devices. However, FRS faces inherent limitations such as data heterogeneity and scarcity, due to the privacy requirements of FL and the typical data sparsity issues of RSs. Models like ChatGPT are empowered by the concept of transfer learning and self-supervised learning, so they can be easily applied to the downstream tasks after fine-tuning or prompting. These models, so-called Foundation Models (FM), fouce on understanding the human’s intent and perform following their designed roles in the specific tasks, which are widely recognized for producing high-quality content in the image and language domains. Thus, the achievements of FMs inspire the design of FRS and suggest a promising research direction: integrating foundation models to address the above limitations. In this study, we conduct a comprehensive review of FRSs with FMs. Specifically, we: 1) summarise the common approaches of current FRSs and FMs; 2) review the challenges posed by FRSs and FMs; 3) discuss potential future research directions; and 4) introduce some common benchmarks and evaluation metrics in the FRS field. We hope that this position paper provides the necessary background and guidance to explore this interesting and emerging topic.
Index Terms:
Federated Recommendation System, Foundation Model, Privacy Preserving, Security,
I Introduction
In the digital age, the exponential growth of information has created a need for systems that navigate, filter, and personalize data for individual users. Recommendation Systems (RS) nowadays become crucial tools for filtering online information and helping users discover products, content, and services that align with their preferences [1]. However, the systems traditionally rely heavily on centralized data collection and processing, posing significant privacy risks and operational bottlenecks. The importance of user privacy has never been greater, particularly with stringent data protection regulations such as the General Data Protection Regulation (GDPR) [2] in Europe, which emphasises the need to store user data on theirs local devices, instead of uploading it to central servers. As a novel approach to address these privacy concerns, Google introduced Federated Learning (FL) [3] as a framework designed to train models across decentralized devices, while keeping data localized. This paradigm shift in data processing leverages the computational capabilities of individual devices for local data analysis. Specifically, FL alternates between local model training at the client and global parameter aggregation from these models on a central server. The integration of FL with RS becomes essential for safeguarding user privacy in recommendation services, which has given rise to the burgeoning field of Federated Recommendation Systems (FRS) [4]. In this field, typically each client consists of a single user’s device. Therefore, unless specified otherwise in this paper, the terms ’user,’ ’client,’ and ’device’ all refer to an individual user. FRS recently have shown promising results in many areas, such as service providing [5, 6], daily scheduling [7], driving planing [8, 9] and more, significantly impacting different facets of daily life. Similar to FL, FRS is required to keep user data local to protect user privacy. However, in most cases, each client typically contains only the data of one user’s accessed item, which is extremely small compared to the total number of item set, creating a serious data scarcity problem. In addition, different users have different behaviours and preferences, which can lead to data heterogeneity. The presence of both issues can lead to sub-optimal models and reduced effectiveness.
Figure 1: The foundation model can integrate information contained in data from various modalities during pre-training. The model can then be adapted for a variety of downstream tasks through adapters such as prompting or fine-tuning.
More recently, the emergence of a novel paradigm for the construction of artificial intelligence (AI) systems has garnered considerable interest in the wake of the remarkable success of ChatGPT [10] and diffusion models [11, 12, 13] in the tasks of language understanding [14, 15, 16, 17] and image generation [12, 13], which we refer to as Foundation Models (FM) [18]. As shown in Fig. 1, this paradigm is built by using self-supervised optimization of training goals to determine how to update model parameters based on model predictions on the large amount of unlabelled training data. This process is referred to as pre-training. Then, FM can utilize a technique, called Adaptation, to turn the general FM into a task-specific one. This process is termed as Fine-tuning or Prompting. Language models, e.g. BERT [19]and RoBERTa [20], are usually trained using the next token prediction goal, which refers to the extent to which the model is able to predict the next token in the sequence. One of the most successful examples of language models is ChatGPT, which is based on GPT-3.5. By training on large amounts of text data, it aligns the capabilities of large language models (LLM) with human intent [18]. Vision models like ViT [21] are typically trained using either contrast learning or diffusion training targets. For contrastive learning [22], images are randomly augmented before evaluating the similarity of the model representations. For diffusion models [11, 12, 13], noise is added to the images and the model is gradually de-noised by the target. There are also multi-modal training targets, some of which separate images and text during training, while others consider both. The foundation model, once trained, is used as a plug-in in combination with adaptations to achieve results for a wide variety of downstream tasks. CLIP [23] and DALL·E [24] are two multi-modal AI models developed by OpenAI. CLIP focuses on understanding images through natural language, while DALL·E focuses on generating images based on text descriptions. While both models are trained on large datasets of images and text, CLIP is primarily used for image retrieval and classification, while DALL·E excels at generating new images that match text descriptions [25].
As FM is frequently trained using a substantial quantity of data from multiple sources, they are capable of incorporating a considerable amount of additional knowledge when performing a specific downstream task. This feature enables FM to provide diverse training data for task-specific models in downstream task scenarios, thus effectively alleviating the problem of data scarcity. Given these considerations, the application of FM to FRS is not only a viable approach but also holds significant research potential. Such integration will drive innovation in FRS. Although the application of FM to FRS has the potential to be highly beneficial, it must be acknowledged that this field is still in its nascent stages, with an insufficient understanding of the challenges, viable methods, and directions for development. This paper aims to bridge this knowledge gap through an in-depth analysis of the integration of FM and FRS. The article provides a comprehensive examination of the motivations and challenges associated with integrating these two paradigms, with a particular focus on several representative technologies. Additionally, it outlines future development trends and their applications. By elucidating the intersection of FRS and FM, this study aims to promote further exploration and innovation within this emerging field, thereby facilitating its rapid advancement.
II Related Work and Contribution
In the field of FRSs, various studies have converged on methodologies, privacy preservation, and challenges, albeit with distinct focal points. Yang et al. [26] discuss the practical implementation and evaluation of FRS, with a focus on system architecture and algorithmic efficiency. Alamgir et al. [27] provides a comprehensive overview of FRS, highlighting techniques, prevailing challenges, and future directions. Javeed et al. [28] focus on the security and privacy concerns in personalized recommendation systems and propose targeted solutions to these challenges. Finally, Sun et al. [29] conduct a survey that compares existing federated recommendation research, highlighting the strengths and limitations of various approaches. All studies emphasize the critical importance of privacy protection and the challenges posed by data heterogeneity and model aggregation. This body of work enhances our understanding of FRS and provides a foundation for future research to address their inherent challenges. Although a number of study focusing on combining FM with FL [30, 31, 32, 33, 34, 35, 36] or RS [37, 38] already exist, our paper is the first work to explore the integration of FM into FRS.
Contribution. This survey aims to provide a clear theoretical framework for applying FM to FRS and to elaborate on their principles and methods of application. By analyzing existing technologies, this paper strives to drive innovation in integrating the pre-training capabilities of FM into FRS. The article discusses in detail the technical challenges and practical issues faced when integrating FM, such as privacy protection, data heterogeneity, communication efficiency, and model generalization capabilities. It also identifies current research gaps and future directions, aiming to guide subsequent academic research and technological development. Additionally, through application studies, this paper demonstrates how to apply the integration of FM with FRS in practical scenarios, thereby deepening the integration of theory and practice.
III Background and Preliminary
III-A Federated Recommendation System
Figure 2: A typical framework for FRS.
(a) Horizontal FRS
(b) Vertical FRS
(c) Transfer FRS
Figure 3: Typical data categories of FRS.
Definition 1 (Federated Recommendation System)
A Federated Recommendation System is a technology that uses distributed algorithms for personalized information filtering. It aims to improve the accuracy of information filtering and the effectiveness of personalized services, while maintaining user privacy.
FRS facilitate collaborative learning by pooling analysis and learning capabilities across multiple clients (e.g. users or locations), without the need for direct exchange of raw user data. Fig. 2 illustrates that this type of system typically involves three processes: 1) client model update, 2) communication, and 3) global aggregation. The client model update allows users to train their models locally on their devices using their own data and then upload the updated intermediate parameters, e.g. the model gradients or parameters can be shared, to the server. The server then performs global aggregation on the parameters sent by all participants, integrating the unique information from each of them. The aggregated parameters are then distributed to the next round of participants, initiating a new training round. The process of uploading and distributing is collectively referred to as communication. These approaches effectively preserves privacy and security.
III-A1 Taxonomy of FRS
Figure 4: The diagram illustrates the taxonomy of FRS based on two main criteria: Data Distribution and Communication Architecture.
As illustrated in Fig. 4, FRS can be categorized based on different criteria [26, 29]. According to used data distribution, FRS can be divided into Horizontal FRS, Vertical FRS, and Transfer FRS. Each type addresses different data collaboration and learning scenarios. Specifically:
  • Horizontal FRS deals with situations where the feature spaces are similar across different entities, but the sample spaces differ, as shown in Fig. 3a. For instance, two customers who shop at the same store may purchase some of the same products. The horizontal method aggregates model updates from various sources to train recommendation models without sharing raw user data. This approach enhances the accuracy and efficiency of recommendation systems while safeguarding user privacy. Horizontal FRS is currently the most common type, with many studies [39, 40, 41] based on the assumption of this data distribution.
  • Vertical FRS is applied when the feature spaces differ, but the sample spaces are similar, as shown in Fig. 3b. This scenario is common when different parties possess different types of data about the same users. An example would be a bank and an online retailer holding distinct perspectives on the same customers, i.e., the bank has credit history while the retailer has shopping history. One challenge is that how to align the identities of customers across different parties. Vertical FRS [42, 43, 44, 45] trains recommendation models by securely integrating different data features between parties, utilizing richer user information for more accurate personalized recommendations and ensuring data privacy and security.
  • Transfer FRS utilizes principles of transfer learning [46] to transfer knowledge from one domain to another, as shown in Fig. 3c. This type is suited for scenarios with significant differences in both feature and sample spaces across parties. Transfer learning allows the system to leverage data and knowledge from the source domain to enhance recommendation performance in the target domain, even when the target domain lacks sufficient data for independent model training. Transfer FRS [47] is particularly beneficial for emerging markets or user groups with limited data, drawing insights from related domains.
(a) Centralized FRS
(b) Semi-decentralized FRS
(c) Decentralized FRS
Figure 5: Typical architecture categories of FRS.
As shown in Fig. 5, FRS can be categorized into centralized, semi-decentralized and decentralized based on their communication architecture. Each type addresses privacy and scalability in different ways:
  • Centralized FRS operates within a FL framework where a central server orchestrates the learning process as shown in Fig. 5a. Clients locally compute updates based on their data and send these updates to the server. The server aggregates these updates to prompt essential information sharing, which is then distributed back to the users. It effectively addresses privacy concerns by allowing the model to learn from decentralized data sources without centralizing the data itself. While this approach improves privacy by not requiring the sharing of raw data, it still relies on a central authority to manage the model. This architecture is typically used to overcome data silo issues [48] and improve the performance of RS without compromising user privacy and data security. Due to its simplicity and intuitive nature, this architecture has become the dominant framework within the field of FRS. There is a great deal of work [49, 39, 40, 41, 47] based on it currently.
  • Semi-decentralized FRS shown in Fig. 5b introduces an intermediate layer between the central server and the users, such as edge servers or devices that can perform additional computational or storage tasks. This setup aims to reduce the communication overhead and latency associated with sending updates to a central server, especially in large-scale applications. A specific example is the SemiDFEGL framework [50], which improves scalability and reduces communication costs by introducing new device-to-device collaborations. It augments local subgraphs with predicted interacted item nodes to exploit high-order collaborative information between users and items in a privacy-preserving manner, which is particularly useful for recommendations based on collaborative filtering and graph neural networks.
  • Decentralized FRS, which employs peer-to-peer communication architecture [51], distributes the learning process completely to all participating devices, without the need for a central server for model aggregation, as shown in Fig. 5c. This approach maximizes privacy and data ownership, but presents challenges in coordinating model updates and ensuring model convergence. Zheng et al. [52] proposed a FRS model called DGREC by adopting a decentralized graph neural network, which constructs a local intra-item hypergraph and a global inter-user graph for each user, allowing users to freely choose whether to disclose their interactions. Li et al. [53] introduced DFedRec, which uses a privacy-aware client-level structured graph to share model parameters only with relevant neighboring users, thereby reducing communication costs and protecting user privacy.
Each of these architectures offers a trade-off between privacy, communication efficiency and the degree of decentralization. Centralized systems simplify model aggregation, but rely on a central authority. Semi-decentralized systems aim to balance efficiency and privacy with an intermediate layer, while decentralized systems offer the highest level of privacy at the cost of more complex model coordination.
III-B Foundation Model
The rapid increase in the performance of computer hardwares, e.g., GPUs, the increasing maturity of transformer architectures, and the public availability of large amounts of training data have been three key factors in the emergence of FM [18]. According to the work of Stanford HAI [18], we have the following definition of FM:
Definition 2 (Foundation Model)
A foundation model is defined as any model trained on extensive data (typically using large-scale self-supervised learning) that is capable of adapting to a wide range of downstream tasks, for instance, through fine-tuning or prompting.
In recent years, the scale and scope of FM have greatly expanded our imagination of potential applications. Such models typically have billions or even trillions of parameters, allowing them to learn more complex patterns and knowledge. They can be adapted to new tasks through fine-tuning or prompting [54, 55]. For instance, the GPT-3 model [56], comprising 175 billion parameters, is capable of adapting and perform various tasks with the aid of natural language prompts, despite the fact that a significant proportion of these tasks have not been explicitly trained. FM is distinguished by two key characteristics [18]: 1) Emergence, which refers to the implicit induction of system behavior from examples, as opposed to explicit design; and 2) Homogenization, which indicates that the method of constructing machine learning systems tends to be unified across a wide range of applications. Although FM is based on self-supervised learning and transfer learning, their scale brings new emergent abilities, and their effectiveness in many tasks has motivated homogenization, e.g., many FMs have evolved based on transformer architectures [19, 20, 21, 56]. While homogenization provides strong leverage, it also requires caution, as defects in foundation models can be inherited by all downstream adapted models. Despite the broad deployment prospects of foundation models, due to their emergent properties, our understanding of how they work, when they fail, and what they can actually do is still quite limited [18], but there has been quite a bit of work [57, 58, 59] that attempts to explain this phenomenon. Moreover, OpenAI proposes a Scaling Law for LLM [60], suggesting that the performance of a model is governed by a power-law relationship involving computational power, model parameters, and data size. This law also applies to downstream tasks [61]. Therefore, understanding and leveraging this power-law relationship when using FM can aid in more efficiently allocating resources and optimizing model performance.
Figure 6: The diagram illustrates the taxonomy of FM based on two main criteria: Data Type and Functionality.
III-B1 Taxonomy of FM
As shown in Fig. 6, similar to FRS, FM can also be classified based on data type or functionality. Specifically, based on the type of data used during training, FM is mainly divided into followings [18]:
  • Language FM: These types of FM mainly deal with textual data and are trained on large textual datasets to understand and generate language [19, 20]. They are very effective in natural language processing (NLP) tasks such as machine translation [62], search [63, 64], and sentiment analysis [65]. Typical language models include BERT [19], GPT-3 [56], and LLama [66], which typically use deep transformer architectures that are able to capture complex linguistic regularities and demonstrate excellent performance on multilingual tasks.
  • Vision FM: Vision FM focuses on processing and understanding image data. These models are able to perform tasks such as object recognition [21, 67, 68], image segmentation [69, 70, 71], and visual reasoning [72, 73] by learning large amounts of image data. For example, DINOv2 [74] and SAM [69] are basic models trained specifically for visual tasks, and using self-supervised learning methods, these models learn valid visual representations without labelling the data.
  • Multi-modal FM: Multi-modal models can simultaneously process and understand multiple types of data, such as text and images. Such models can excel in cross-modal tasks such as image captioning and visual quizzing [75, 76, 77] by integrating information from different modalities. CLIP [23] and DALL·E [24] are representative of such models that are able to understand the relationship between an image and its corresponding textual descriptions, demonstrating flexibility and robustness in handling multiple data types.
Functionally, FMs are generally categorized into two types [18, 78, 38]:
  • Discriminative FM: The main task of discriminative FM is to distinguish or predict specific outputs from given input data. These models are generally based on the BERT family and are more concerned with learning decision boundaries from the data to perform tasks like classification [79, 80] and regression [81].
  • Generative FM: The core goal of generative FM is to learn the distribution of the data and be able to generate new data samples [25]. These models, such as GPT-3, DALL·E, etc., are able to capture the underlying structure of the data and thus generate new instances that are similar but different from the training data. Generative FM has a wide range of applications in areas such as the generation of text [82] and images [12, 13].
III-B2 Adaptation
Fig. 1 illustrates a commonly used technique for pre-trained models, known as the Adaptation. This technique adds new lightweight layers while maintaining the original parameters of the pre-trained model, enabling fine-tuning and extension for specific tasks [83, 41]. This approach is suitable for multi-task learning and tasks performed in resource-limited environments, such as few-shot learning [84]. Adaptations have found wide applications in natural language processing and computer vision tasks [85, 86, 87] . According to [84], adaptation methods are categorized into three main types:
  • Prompt-based Fine-tuning Adaptation Methods are inspired by the principles of prompt learning in NLP, adapts the model to a specific downstream task by using a portion of the fixed text input as a learnable vector and fine-tuning it using downstream task data. For instance, CoOp [88] introduces the concept of prompt learning into the adaptation of downstream tasks for multimodal pre-trained foundation models by automatically constructing contextual prompts using learnable word embeddings. CoCoOp [89] builds upon this by creating a meta-network that learns image features, which are then combined with prompt vectors to enhance the generalization performance of CoOp on new category data.
  • Adapter-based Fine-tuning Adaptation Methods entail the incorporation of small, trainable modules called Adapters into a pre-trained model. These adapters are subjected to fine-tuning, rather than the entire model, thereby enabling the model to adapt to novel tasks with minimal alterations to its original parameters. CLIP-Adapter [87] adds simple adapter structures within the CLIP model that are fine-tuned using a small dataset. This method allows the model to adapt to new tasks by learning task-specific features without the need for extensive retraining of the entire model. CALIP [90] enhances the cross-modal interaction capabilities of adapters by integrating text and image features using attention maps and incorporating two trainable linear layers before and after this integration. LoRA [85] adjusts the weights of a pre-trained model by multiplying the weight matrices with a low-rank matrix, allowing for fine-tuning of the model without significantly increasing the number of parameters.
  • External Knowledge-based Adaptation Methods integrate external knowledge (e.g., from knowledge graphs or additional language model queries) into the adaptation process. This is particularly advantageous for enhancing the model’s comprehension and efficacy in tasks where domain-specific knowledge is vital but data is scarce. CuPL [91]enhances zero-shot classification performance by asking questions to a LLM and generating multiple descriptive statements for each category, enriching the semantics of the categories. LaBo [92] uses GPT-3 to generate a large pool of candidate feature descriptors. It then selects the optimal subset of descriptors for classification decisions, considering the distinctiveness of features for other categories and the coverage for the current category.
IV Federated Recommendation Systems with Foundation Models
As described earlier in Fig. 2, a typical FRS typically consists of three stages [27]: client model update, communication, and global aggregation. Integration with FM should also occur at these three stages.
IV-A Client Model Update
In FRS, clients have the following characteristics [26, 27]:
  • As mentioned earlier, a major advantage of FRS is privacy. The data required for model training in FRS, such as user information and interaction history, is privacy sensitive and therefore must remain on the user’s device, creating data silos. This approach not only complies with data privacy regulations (such as GDPR), but also increases user trust. However, this means that each user has a relatively small amount of data compared to the entire dataset, and each user only accesses a portion of the item set, leading to data sparsity.
  • Each client independently trains the model on its local data. This allows the model to better adapt to each user’s unique preferences and behaviors, improving personalization and accuracy of recommendations. The resulting data distribution often does not satisfy the independent and identically distributed (IID) assumption, leading to the problem of data heterogeneity.
  • In addition, user devices are typically consumer products such as mobile phones and personal computers, which are characterized by unstable communication and limited computational resources. This requires keeping the computational load of the models deployed on the client and the amount of information exchanged with the server as low as possible.
FM is typically pre-trained on large, diverse datasets, providing it with rich knowledge and strong representational capabilities. This pre-training provides it with prior knowledge that allows it to quickly adapt to specific customer data through fine-tuning [18]. As a result, FM can learn general feature representations that provide a degree of adaptability to different data distributions. Applying these models during the client update phase in FRS can significantly improve the performance of local models. Specifically:
  • FM can apply knowledge learned from large corpora to local data through transfer learning. For example, a pre-trained FM can be effectively fine-tuned on a small amount of local data in the client to adapt to specific user behavior data, thereby achieving good performance in specific downstream tasks such as providing more accurate recommendations [38]. In addition, by fine-tuning the base model locally, sensitive data does not need to leave the device, protecting user privacy [93]. As a result, customers can achieve better model performance with limited data.
  • FM can also capture complex user preferences and behaviors. For example, FM has strong semantic understanding capabilities that allow them to better understand users’ search queries [63, 64], comments [37], and other textual data [94] to provide recommendations that better meet user needs. In addition, as user behavior changes, FM can quickly adapt and adjust recommendation strategies to ensure that recommendation results remain consistent with the user’s current interests and needs. By applying FM during the client update phase, a higher level of personalized recommendation can be achieved.
However, it is important to note that when FM is applied to data that differs significantly from the training distribution, performance degradation can occur. This problem, known as out-of-distribution generalization [93], is a challenge that FM must overcome. In addition, if there are biases in the training data set, FM may learn and reinforce these biases, leading to inequitable results across different data distributions [95, 96, 97]. Although fine-tuning FM requires significantly less resources than training from scratch, it still typically requires significant computational resources for effective fine-tuning and updating [18, 36]. This requirement may limit its use on resource-constrained clients, especially in FRS where each user represents a client, potentially limiting the use of FM due to limited computational resources.
IV-B Communication
In FRS, communication between clients and the server is a critical phase. The primary task at this stage is to transmit model update information from each client to the central server while ensuring data security and communication efficiency during the transmission process. In traditional FRS frameworks, communication between clients and the server typically involves transmitting large amounts of update information, which may include model parameters, gradients, or other relevant statistical data. To ensure efficient and secure communication, common strategies include data compression [98, 99, 100], encrypted transmission [101, 102], and secure aggregation [103, 104, 105]. However, with the increasing data scale and higher privacy protection requirements, these traditional methods face significant challenges [36].
FMs, especially LLMs, possess strong representation capabilities, enabling them to extract key features from model updates through efficient data representation [55, 106] and semantic compression techniques [107, 108]. By removing redundant information and unnecessary data, and retaining only essential information, FM can transform this into a more compact representation, significantly reducing the data transmission volume between clients and the server. Optimizing communication protocols [109] and enhancing data compression and security during this process can markedly improve the overall performance of the system.
IV-C Global Aggregation
In FRS, the global aggregation phase of the server is responsible for combining model updates from all clients into a single global model. This phase is of critical importance in ensuring the overall performance and effectiveness of personalized recommendations. Server global aggregation is characterized by the following key features [110]: First, server global aggregation must handle updates from a large number of clients that contain different data and user behaviors. To ensure the overall quality of the model and the accuracy of the recommendations, the server must effectively integrate this diverse data. Second, the global aggregation process must be sufficiently robust to handle potentially noisy data and anomalous updates. Finally, the scalability of the server’s global aggregation is critical, especially when dealing with a large number of clients and massive amounts of data, requiring the system to efficiently process and aggregate updates. In each iteration, given resource constraints and privacy concerns, the central server may only communicate with a subset of clients. Researches [111, 112] indicate that clients should not participate in successive training rounds, as attackers could potentially infer information about the client from transmitted gradients or model parameters. The client selection strategy determines which clients will participate in the current training round based on factors such as computational power, network stability, data diversity, and data quality [113]. For example, client selection can also be done by random sampling or based on statistical characteristics of the data to ensure that the model learns from diverse data sources, thereby increasing its generalization ability.
Context-aware aggregation represents a critical application of FM in global server aggregation. Traditional aggregation methods, such as naive weighted averaging proposed in FedAvg [3], may not fully exploit the contextual information present in the data. FM, however, is capable of intelligently aggregating updates by understanding the semantics and contextual relationships of each participant’s updates [70]. To illustrate, when FM receives customer updates, it analyzes the contextual information of each update, such as user behavior patterns, preferences, and history. This information enables FM to better comprehend the significance and value of each update and to assign distinct weights accordingly [32]. This context-aware aggregation method can more accurately reflect user preferences and behaviors, thereby enhancing the overall performance of the model. Dynamic weighting permits the system to adapt more effectively to users’ personal needs and to provide recommendations that are more aligned with their interests.
With its extensive knowledge base [114], FM can utilize this knowledge to make more informed decisions during the aggregation process. Knowledge-based aggregation strategies can capitalize on FM’s knowledge and reasoning capabilities to enhance the efficacy of aggregation. To illustrate, in the event that certain customers lack sufficient data, FM can leverage pertinent information from its knowledge base to infer relationships and patterns between data, thereby supplementing and improving aggregation. The key concept of Mixture of Experts (MoE) [115] is that different expert models specialize in different parts of a problem. By appropriately weighting and combining their outputs, better overall performance can be achieved. The MoE architecture has a similar function to aggregation strategies in FRS. Therefore, applying this architecture to the aggregation strategy in FRS may result in improved recommendation performance.
In terms of anomaly detection, FM also performs exceptionally well [116, 117]. In FRS, client updates may be influenced by various factors, leading to noisy or abnormal data [29]. If these abnormal updates are directly used in global aggregation, they may affect the overall quality of the model. FM can utilize its strong anomaly detection capabilities to analyze client updates, identify and handle these anomalies, ensuring the robustness of global aggregation. For instance, if a client’s data update exhibits a notable divergence from the majority of client updates, it may be regarded as noise or an anomaly. FM has the capacity to elect to disregard or assign a diminished weight to it, thus reducing its deleterious impact on the global model and ensuring the stability and robustness of the global model. Liu et al. [37] observed in their study that FM is capable of effectively filtering out noisy data by identifying aberrant patterns, thereby enhancing the stability and reliability of the aggregation process.
This application direction has considerable potential for future development and is worthy of further research and investigation.
V Challenges
The application of FM to FRS marks the advent of a new frontier in personalized content recommendation. This approach promises to further enhance user experience by employing more powerful models than traditional deep models while protecting user privacy. However, this promising combination comes with many challenges, involving various aspects such as data privacy and security, data sparsity and sample imbalance, model training and optimization, model interpretability and transparency, and so on. It is therefore of the utmost importance to gain a comprehensive understanding of these issues and to develop effective strategies to address them. This is essential if FRS is to be successfully deployed and operated in a secure yet powerful manner. In this section, we will examine these challenges in detail, explore their implications and discuss potential strategies to overcome them.
V-A Data-Related Challenges
  • Data Privacy: In a federated setting, the protection of user privacy is of paramount importance. The model must ensure that training and recommendations are performed without disclosing personal data. However, FMs, such as GPT-3, have the potential to memorize and reproduce data from the training set, thereby potentially leaking sensitive information. Additionally, if the generated data is too similar to the original data, there is a risk of privacy infringement. This may involve the use of additional encryption techniques to safeguard user data. Differential privacy [118] protects individual data by adding noise during model training, thus reducing the risk of data leakage while maintaining model performance. Homomorphic encryption [119] allows computations on encrypted data, enabling model training without decrypting the data, effectively preventing data theft during transmission and processing. Furthermore, machine unlearning techniques [120] can remove specific user data from the foundational model, ensuring that these data no longer influence the model’s predictions and outputs. This approach is particularly effective in the context of user deletion requests or in the context of compliance with privacy regulations such as the GDPR.
  • Data Security: The integration of FRS and FM also necessitates the equal consideration of data security. Participants may be subjected to malicious attacks, including Member Inference Attacks [121], Data Reconstruction Attacks [122], and Poisoning Attacks [123], which could compromise the model or alter the data. Consequently, ensuring the integrity and confidentiality of data during transmission and storage represents a significant challenge. The application of secure multi-party computation (MPC) [124] can distribute computational tasks among multiple participants, ensuring that no single party can obtain complete data, thereby enhancing security. Additionally, blockchain [125, 126] can be used to record data access and operation logs, ensuring data integrity and traceability. This also helps improve the transparency and security of the system.
  • Data Sparsity and Scarsity: RS frequently encounters the challenge of data sparsity, where there is limited interaction data between users and items. This leads to decreased model performance. This issue is particularly pronounced in a federated setting, where each participant has limited data. In specific domains or tasks, insufficient data can cause severe data scarcity, further restricting model training and performance. Although FM has the ability to perform zero-shot and few-shot learning, they still need optimization to handle extremely sparse data. It is therefore of great importance to utilise data augmentation techniques [127, 128] in order to generate a greater quantity and variety of training samples, thereby increasing the overall data set. Alternatively, knowledge transfer [129] between different domains can be employed in order to leverage the rich data sets from other areas, thus alleviating the issue of data sparsity and improving the overall model performance in the current domain. Despite the ability of generative FM [130, 131] to produce a vast quantity of synthetic data, the quality of this data may not be guaranteed. If the FM utilises low-quality or biased data sources, the generated data may inherit these issues and even exacerbate existing biases and errors [132]. Wang et al. [22] found that generated data is not always beneficial for representation learning. For example, in contrastive learning, the optimal performance was achieved when the mix ratio of real to generated data was approximately 10:1. Consequently, when generating synthetic data, it is of paramount importance to establish quality standards and utilise evaluation metrics to assess data quality. Furthermore, it is essential to implement quality control mechanisms to filter out low-quality or biased data.
  • Sample Imbalance: In a federated setting, there are significant differences in data size and distribution among different clients [133, 134, 135]. Similarly, in RS, item labels and user behaviors often exhibit a long-tail distribution [136]. These issues are referred to as sample imbalance, where part of classes have much more samples than others in the dataset. This can cause the model to overfit the majority classes while having poor generalization for minority classes, thus affecting the model’s training effectiveness. To address this, additional methods are required to balance these differences and ensure the model’s fairness and accuracy. Resampling techniques typically involve either oversampling or undersampling to balance the dataset [137]. Oversampling increases the number of minority class samples, while undersampling reduces the number of majority class samples. Weighted loss functions can assign higher weights to minority class samples during model training, enhancing the model’s ability to recognize minority classes. This approach can enhance the fairness and accuracy of the model without modifying the data distribution.
V-B Model-Related Challenges
  • Model Synchronization: In a federated setting, models from various participants frequently communicate with the server for synchronization, resulting in high communication costs and complex system management issues. This is particularly problematic for large FM due to their vast number of parameters [18], which require significant computational resources and time for transmission and synchronization. The use of gradient compression techniques [138] to reduce the amount of data transmitted during model synchronization can effectively lower communication costs and enhance synchronization efficiency. Furthermore, the implementation of asynchronous update strategies [139], which permit participants to independently update their models and periodically perform global synchronization, can reduce waiting times during synchronization and enhance the overall efficiency of the system.
  • Model Heterogenity: Model heterogeneity [140] is typically attributable to a multitude of factors, including disparate types, sizes, and architectures of models among various clients. This can be attributed to differences in client hardware, data distribution, or business needs. In certain instances, different clients may utilize distinct model structures. For instance, some clients may employ simple linear models, while others may utilize complex deep learning models [141]. This renders it challenging to deploy or fine-tune a unified foundation model in a federated setting. The design of adaptive training algorithms [142, 143] enables the dynamic adjustment of model parameters based on the computational capabilities and data characteristics of each participant, thereby enhancing the model’s adaptability and performance in heterogeneous environments. Furthermore, knowledge transfer [47] among different participants facilitates the sharing of insights and the mitigation of data and computational resource disparities, resulting in an improvement in the overall performance of the model.
  • Model Interpretability: FM is often represented by black-box models, which makes it challenging to comprehend their internal workings. This lack of interpretability can give rise to issues of trust and compliance in practical applications. In FRS, model interpretability is of paramount importance for establishing user trust and meeting regulatory requirements [144]. Therefore, it is vital to employ explainable AI techniques [145] to provide transparent and understandable explanations for the decisions made by FM. For instance, attention mechanisms [146] and feature importance analysis [147] can be utilized to elucidate model decisions. Furthermore, the utilisation of generative FM [148] to generate natural language explanations for recommendations can provide users with the context and rationale behind the model’s decisions, thereby enhancing user trust and acceptance of the model.
  • Model Transparency: The concept of model transparency entails the open and auditable process of model decision-making. In the context of FRS, transparency plays a pivotal role in fostering user trust, necessitating the assurance of model fairness and non-discrimination. This encompasses the monitoring of model training processes, the validation of model updates, and the tracing of generated data. A lack of transparency can diminish the credibility of the model and increase compliance risks. FRS with FM should have transparent training protocols [149] that record and monitor each step of the model training process. This ensures transparency, thereby enhancing system credibility and compliance. Similarly, establishing a detailed logging system [48] to record every model update and operation ensures traceability, helping to identify and correct issues within the model and improve overall system transparency.
VI Future Directions
The emerging field of applying FM to FRS presents both opportunities and challenges. The combination of the powerful representation capabilities of FM and the data privacy protection advantages of FRS can enhance the performance of recommendation services and address some of the challenges that traditional FRS methods struggle to overcome. This section discusses some key future directions for FRS with FM, aiming to provide valuable guidance for future research and applications. These directions can offer higher quality personalized recommendation services while protecting user privacy. As technology continues to advance, it is reasonable to posit that FM will play an increasingly important role in future FRS.
VI-A Data Augmentation
FMs, especially generative FMs such as LLMs and DMs, have the ability to generate training data needed for downstream tasks such as text and image generation [150, 151]. They can be used to generate more user interaction data, alleviating data scarcity problems. In FRS, data from different participants may be unbalanced or insufficient. Using FM for data augmentation can effectively increase the variety and amount of training data.
Through data augmentation, FM can help build more comprehensive user profiles and item characteristics, thereby improving the personalization and effectiveness of RS. For example, virtual interaction records of users or additional item descriptions can be generated to better capture user interests and preferences [152]. However, as mentioned earlier, the generated data may have quality issues that could degrade the performance of the RS and introduce biases and errors. Therefore, it is critical to perform rigorous validation and cleaning of the generated data to ensure that the synthetic data is statistically consistent with the real data and to reduce the noise introduced by the synthetic data [132]. Similarly, FM should be combined with privacy-preserving techniques, such as differential privacy, to ensure that no sensitive user information is revealed during the generation of synthetic data. In addition, FM should be used to generate diverse data to cover different scenarios and user behaviors, reducing the risk of bias and over-fitting [153].
VI-B Cold-Start Recommendation
The core of the cold start problem is data sparsity, meaning that new users and new items lack sufficient historical interaction data [154]. This makes it difficult for the model to capture users’ true interests and preferences, which affects the accuracy and effectiveness of recommendations. Because FM has been pre-trained on large amounts of textual data, it has learned rich semantic information that helps generate high-quality representations of users and items. Therefore, FM has strong zero-shot and few-shot learning capabilities, allowing it to provide effective recommendations for cold start users and items even without sufficient interaction data [155, 41]. In addition, FM can transfer knowledge learned in one domain to another through techniques such as fine-tuning and prompt learning. This means that even in cold start situations, the model can use previously learned knowledge to make recommendations, thereby improving the performance of the recommendation system.
It is important to note that in a federated environment, collecting and using enough information to solve the cold start problem while protecting user privacy is a significant challenge. Ensuring the privacy and security of user data is a fundamental requirement for any FRS. In addition, multi-task learning [156] can improve the adaptability and generalization of basic models across different tasks and domains. For example, training on both recommendation and classification tasks simultaneously enables the model to make effective recommendations in more scenarios.
VI-C Multi-Modal Recommendation
FM has the ability to process and understand different modalities of data, such as text, images, audio, and video [18]. In FRS, the integration of multi-modal data enables FM to build more comprehensive and multi-dimensional user profiles, thereby improving the personalization of recommendations [157, 158]. These rich user profiles can more accurately reflect the user’s interests and preferences, helping to deliver recommendations that better meet the user’s needs.
It is important to note that a major challenge of multi-modal data is its heterogeneity, meaning that different modalities of data have different characteristics and structures [159]. For example, text data is sequential, while image data consists of two-dimensional pixels. This heterogeneity increases the complexity of data fusion and processing. Therefore, the development of unified data representation methods that map different modalities of data into a common representation space is essential for model processing and integration.
In a federated environment, special attention must be paid to the privacy of multi-modal data. Multi-modal data may contain more sensitive information, such as the user’s voice or video, which requires more stringent privacy measures to effectively handle and protect this data. In addition, the processing and integration of multi-modal data requires significant computational resources and time. The training and inference processes of FM on multi-modal data can be more complex and time-consuming, making it critical to improve computational efficiency.
VI-D Real-Time Recommendations
Real-time recommendation is a critical task in FRS, which aim to dynamically deliver personalized content based on real-time user behavior and contextual data [160]. This recommendation model emphasizes immediate processing and response to data, ensuring that users receive the most relevant recommendations in the shortest possible time. FM excels at understanding contextual information, which enhances their performance in processing user queries, item descriptions, and other textual data to improve recommendation accuracy and relevance.
However, real-time recommendations require processing and responding to user requests in extremely short time frames, which places high demands on computational resources and system latency. FM typically requires significant computational resources, which can lead to increased latency and negatively impact the user experience. In addition, limitations on the length of the input context can cause performance issues when dealing with long sequences of user behavior or a large number of candidate items.
In real-time recommendations, FM can also reflect biases from the training data in its recommendations, leading to unfair biases. For example, popular items may be preferentially recommended due to their high frequency in the training data, overshadowing long-tail items [136]. Therefore, it is crucial to explore efficient model compression and acceleration techniques, such as distillation [161] and pruning [162], to reduce computational complexity and latency. Introducing strategies such as sliding windows [163] can manage and optimize the input of contextual information, thereby improving the model’s performance in handling long behavioral sequences. Incremental learning methods [164, 165]can also be used to update the FM, and model distillation techniques can reduce model size, thereby lowering communication overhead.
VI-E Enhanced Recommendation Explainability
The federated nature and recommendation characteristics of FRS require transparent explanations of recommendation results to users. This transparency helps users understand the logic and reasoning behind the recommendations, thereby increasing their trust and satisfaction. Language FMs, pre-trained on large amounts of text data, have strong natural language generation capabilities. They can generate clear and coherent textual explanations while referencing relevant background knowledge, helping users to understand the reasons behind recommendations [166, 167]. For example, models such as ChatGPT can generate detailed recommendation explanations that explain why a particular item is suitable for the user [168]. However, generating detailed recommendation explanations often requires significant computational resources [169]. This is especially true for LLMs, which can increase latency and negatively impact the user experience. In addition, these LLMs may exhibit biases present in the pre-training data, resulting in unfair tendencies in the generated explanations, such as gender or race biases. Therefore, it is important to apply model compression and optimization techniques to reduce the computational burden of FM, and to design debiasing methods and fairness constraints to minimize biases in recommendation explanations. Furthermore, it is crucial to incorporate user feedback mechanisms to continuously improve the quality of recommendation explanations. By leveraging user feedback, strategies for generating recommendation explanations can be adapted and optimized, thereby increasing user satisfaction and trust.
VI-F Adavanced Metrics
Traditional evaluation metrics, which will be detailed in VII-B, are designed for specific tasks in recommendation scenarios, such as rating prediction or item ranking. However, the strength of FM in RS lies in its powerful generative capabilities, which allow it to generate items that have never appeared in historical data and recommend them to users. In addition, FM can generate explanations and user interaction content, which requires more complex and comprehensive evaluation criteria. In this context, how to evaluate the recommendation capabilities of these systems remains an open question. Therefore, FRS with FM requires new evaluation metrics tailored to the specific applications of FM in RS, especially in generative recommendation.
VII Resources
VII-A Datasets
Figure 7: Two types of Feedbacks.
TABLE I: Datasets with Explicit Feedback.
Context
Dataset #User #Item #Feedback Sparsity User Item Feedback Timestamp
Amazon Beauty 1,210,271 249,274 2,023,070 99.99%
Amazon Books 8,026,324 2,330,066 22,507,155 99.99%
Amazon CDs 1,578,597 486,360 3,749,004 99.99%
Amazon Cell Phone 2,261,045 319,678 3,447,249 99.99%
Amazon Clothing 3,117,268 1,136,004 5,748,920 99.99%
Amazon Games 826,767 50,210 1,324,753 99.99%
Amazon Garden 714,791 105,984 993,490 99.99%
Amazon Home 2,511,610 410,243 4,253,926 99.99% $\checkmark$ $\checkmark$
Amazon Kindle 1,406,890 430,530 3,205,467 99.99% $\checkmark$ $\checkmark$
Amazon Movies 2,088,620 200,941 4,607,047 99.99% $\checkmark$ $\checkmark$
Amazon Music 478,235 266,414 836,006 99.99% $\checkmark$ $\checkmark$
Amazon Sports 1,990,521 478,898 3,268,695 99.99% $\checkmark$ $\checkmark$
Amazon Toys 1,342,911 327,698 2,252,771 99.99% $\checkmark$ $\checkmark$
Amazon Reviews Amazon Video 426,922 23,965 583,933 99.99% $\checkmark$ $\checkmark$
ML-100K 943 1,682 100,000 93.70% $\checkmark$ $\checkmark$ $\checkmark$
ML-1M 6,040 3,952 10,000,209 95.81% $\checkmark$ $\checkmark$ $\checkmark$
ML-10M 69,878 10,681 10,000,054 98.69% $\checkmark$ $\checkmark$
MovieLens ML-20M 138,493 27,278 20,000,263 99.47% $\checkmark$ $\checkmark$
Yelp2018 1,326,101 174,567 5,261,669 99.99% $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
Yelp2020 1,968,703 209,393 8,021,122 99.99% $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
Yelp2021 2,189,457 160,585 8,635,403 99.99% $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
Yelp2022 1,987,929 150,346 6,990,280 99.99% $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
Yelp Yelp-full 5,556,436 539,254 28,908,240 99.99% $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
Anime 73,515 11,200 7,813,737 99.05%
Book Crossing 105,284 340,557 1,149,780 99.99% $\checkmark$ $\checkmark$
Douban 738,701 28 2,125,056 89.73% $\checkmark$ $\checkmark$
Epinions 116,260 41,269 188,478 99.99% $\checkmark$ $\checkmark$
Goodreads 876,145 2,360,650 228,648,342 99.99% $\checkmark$ $\checkmark$
Jester 73,421 101 4,136,360 44.22%
Netflix 480,189 17,770 100,480,507 98.82% $\checkmark$
Yahoo Music 1,948,882 98,211 11,557,943 99.99% $\checkmark$
TABLE II: Datasets with Implicit Feedback.
Context
Dataset #User #Item #Feedback Sparsity User Item Feedback Timestamp
MIND_large_train 711,223 27,047 83,507,374 99.57% $\checkmark$
MIND_large_dev 255,991 6,998 14,085,557 99.21% $\checkmark$
MIND_small_train 50,001 20,289 5,843,444 99.42% $\checkmark$
MIND MIND_small_dev 50,001 5,370 2,740,998 98.98% $\checkmark$
QK-video 5,022,750 3,753,436 142,321,193 99.99% $\checkmark$ $\checkmark$
QK-article 1,325,838 220,122 46,111,728 99.98% $\checkmark$ $\checkmark$
QB-video 34,240 130,647 1,701,171 99.96% $\checkmark$ $\checkmark$
Tenrec QB-article 24,516 7,355 348,736 99.81% $\checkmark$ $\checkmark$
Adressa 15,514 923 2,717,915 81.02% $\checkmark$ $\checkmark$
Foursquare 1,083 38,333 227,428 99.45% $\checkmark$ $\checkmark$
Gowalla 196,591 950,327 6,442,890 99.99%
Last.FM 1,892 17,632 92,834 99.72% $\checkmark$
Pinterest 55,187 9,911 1,445,622 99.74%
Steam 2,567,538 32,135 7,793,069 99.99% $\checkmark$ $\checkmark$ $\checkmark$
TaFeng 32,266 23,812 817,741 99.89% $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
Tmall 963,923 2,353,207 44,528,127 99.99% $\checkmark$ $\checkmark$
In FRS, the datasets used can be categorized into two types based on user feedback: explicit feedback and implicit feedback, as shown in Fig. 7. Explicit feedback includes direct responses from users about their preferences. This typically includes ratings (such as a 1-5 scale), user comments, and like/dislike statements. Such data is clear and provides straightforward insights into user preferences, making it highly valuable for training recommendation models. Common datasets with explicit feedback include the following shown in Table I:
  • Amazon Reviews [^1] https://cseweb.ucsd.edu/ jmcauley/datasets.html#amazon_reviews: The Amazon Reviews dataset is a large-scale dataset that contains product information across various categories such as Books, CDs, and Music. It includes reviews (ratings, text, helpfulness votes) and product metadata (description, category information, price, brand, and image features). There are three updated versions of this dataset from the years 2014 [170, 171], 2018 [172], and 2023 [173].
  • MovieLens Datasets [^2] https://grouplens.org/datasets/movielens/: The MovieLens datasets [174], initially released in 1998, capture individuals’ stated movie preferences. These preferences are recorded as tuples, with each tuple showing a person’s rating (from 0 to 5 stars) for a movie at a specific time. Users enter these ratings through the MovieLens website, which provides personalized movie suggestions based on these ratings.
  • Yelp Datasets [^3] https://www.yelp.com/dataset: This dataset is a subset of Yelp’s business, review, and user data. It was initially developed for the Yelp Dataset Challenge, which allows students to study or analyze Yelp data and present their insights. In total, there are four versions of the Yelp datasets.
  • Anime [^4] https://www.kaggle.com/datasets/CooperUnion/anime-recommendations-database: This dataset collects user preference data from the MyAnimeList website. It contains information from 73,516 users on 12,294 anime titles. Users can add anime to their completed list and rate them, and this dataset compiles these ratings.
  • Book Crossing [^5] https://grouplens.org/datasets/book-crossing/: The Book-Crossing dataset is a well-structured collection of data collected by Cai-Nicolas Ziegler in a 4-week crawl from the Book-Crossing community. This dataset primarily comprises user interactions that include book ratings, ranging from 0 to 10.
  • Douban [^6] https://www.kaggle.com/datasets/utmhikari/doubanmovieshortcomments: The Douban Movie dataset is a Chinese website where internet users can post their opinions and comments about films. This dataset contains over 2 million short comments on 28 movies from the Douban Movie website.
  • Epinions [^7] https://cseweb.ucsd.edu/ jmcauley/datasets.html#social_data [175]: This dataset was collected from Epinions.com, a popular online consumer review site. It includes trust relationships between users and covers a period from January 2001 to November 2013.
  • Goodreads [^8] https://www.kaggle.com/datasets/jealousleopard/goodreadsbooks: This dataset includes reviews from the book review website Goodreads, along with various attributes describing the books. Importantly, the dataset captures different levels of user interaction, from adding books to a shelf, to rating them, to reading them.
  • Jester [^9] https://eigentaste.berkeley.edu/dataset/: The Jester dataset focuses exclusively on jokes. Users of the Jester online platform rate jokes and these ratings are then used to personalize joke recommendations for them.
  • Netflix [^10] https://www.kaggle.com/datasets/netflix-inc/netflix-prize-data: Netflix provided a training dataset consisting of 100,480,507 ratings from 480,189 users for 17,770 films. Each rating is represented as a set of four elements: ¡user, movie, rating date, rating score¿. The user and the movie are identified by integer IDs, and the rating scores range from 1 to 5 stars, also as integers.
  • Yahoo Music [^11] https://webscope.sandbox.yahoo.com/catalog.php?datatype=r: The Yahoo Music dataset is known for its large scale and diversity. It contains a large collection of user ratings on different musical elements such as tracks, albums, artists and genres. This dataset was used in the KDD-Cup 2011 competition, where participants were asked to analyse user preferences in music based on these ratings.
Implicit feedback, on the other hand, is derived from user actions that indirectly indicate preferences, such as bookmarks, video/music play history, or click-throughs. Although implicit feedback does not directly express user likes or dislikes, it is rich and captures user behaviour more comprehensively.
  • MIND [^12] https://msnews.github.io/ [176]: The MIND dataset, sourced from the Microsoft News website, is a large-scale collection of approximately 160,000 English news articles and over 15 million user interaction records. It has been designed to advance research in news recommendation systems. It includes detailed textual content for each story and anonymized user interaction data to ensure privacy.
  • Tenrec [^13] https://github.com/yuangh-x/2022-NIPS-Tenrec [177]: The Tenrec dataset is a comprehensive benchmark dataset for RSs, featuring user interactions from two recommendation platforms across four dataset files: QK-video and QB-video for video actions, and QK-article and QB-article for article actions.
  • Adressa [^14] https://reclab.idi.ntnu.no/dataset/ [178]: The Adressa dataset is a corpus of Norwegian news articles related to anonymous users. It is a collaborative project between the Norwegian University of Science and Technology and Adressavisen. The objective is to gain insight into the nature of news articles and their readers.
  • Foursquare [^15] https://sites.google.com/site/yangdingqi/home/foursquare-dataset [179]: This dataset comprises check-in data from New York City, collected over a period of approximately ten months (from 12 April 2012 to 16 February 2013). It encompasses 227,428 check-ins in New York City, with each check-in recorded with its timestamp, GPS coordinates, and detailed venue category.
  • Gowalla [^16] https://snap.stanford.edu/data/loc-gowalla.html [180]: Gowalla is a location-based social networking website where users can post their whereabouts by checking in. The dataset comprises data collected from the public API, which represents an undirected friendship network with 196,591 nodes and 950,327 connections. Additionally, it records 6,442,890 check-ins made by these users between February 2009 and October 2010.
  • Last.FM [^17] https://grouplens.org/datasets/hetrec-2011/ [181]: The Last.FM dataset represents a valuable resource that has been extensively utilized in the field of music information retrieval and RSs. It captures detailed information regarding music listening events from users. Each listening event is further enhanced with user demographics and specific descriptors that reflect their music tastes and consumption behaviours.
  • Pinterest [^18] https://github.com/edervishaj/pinterest-recsys-dataset [182]: The Pinterest dataset represents a valuable resource for a variety of research and analytical purposes. It encompasses a diverse range of data, including images, user features, interests, and user interactions.
  • Steam [^19] https://github.com/kang205/SASRec [183]: The Steam dataset is a collection of information about games published on the Steam platform. It includes details such as game names, release dates, genres, developers, publishers, and other relevant information.
  • TaFeng [^20] https://www.kaggle.com/datasets/chiranjivdas09/ta-feng-grocery-dataset: The TaFeng dataset is a comprehensive collection of supermarket shopping data, including detailed transaction records from the Ta Feng supermarket in Taiwan, covering a period from November 2000 to February 2001. The dataset comprises a variety of data points, including customer demographics, product categories, and detailed item descriptions along with quantities purchased.
  • Tmall [^21] https://tianchi.aliyun.com/dataset/53[184]: The Tmall dataset is a comprehensive collection from Tmall, comprising anonymized user shopping records over a six-month period up to and including the ”Double 11” event. It should be noted that the data is selectively sampled to address privacy concerns.
Both types of feedback play a critical role in the development of FRSs, providing diverse insights into user preferences that help improve the accuracy and relevance of the recommendations provided.
VII-B Metrics
In the field of FRSs, the use of evaluation metrics is fundamental to assessing and refining the performance of our algorithms. Metrics serve as a quantitative lens through which we can observe how closely the system’s suggestions match users’ actual interests and preferences.
For predicting how well a system can estimate user preferences, there are several metrices measuring the prediction errors to judge the accuracy of the predictions [185], such as Mean Absolute Error (MAE), Mean Squared Error (MSE) and Root Mean Squared Error (RMSE). When it comes to classifying items, i.e., determining whether a user will like a product or not, we look at metrics such as Precision, Recall, Hit Ratio (HR), F1 Score, Accuracy and AUC. These tell us how correctly the RS is classifying items, and the F1 score helps us balance the Precision and Recall.
Then to measure the item ranking ability, which is about listing recommendations in the right order, Average Precision (AP) and Mean Average Precision (MAP) are key to this, as they assess the quality of the order of recommendations. Metrics such as Mean Reciprocal Rank (MRR) [186], Normalized Mutual Rank (NMR) [187] and Normalised Discounted Cumulative Gain (NDCG) [188] also contribute by assessing how well the top recommended items are ranked.
On a broader scale, we consider recommendation-centric metrics such as Diversity, which ensures that a variety of items are suggested, and Coverage, which measures how many items from the catalogue are recommended. There are also some user-centric metrics include Novelty, which measures how new or surprising the recommendations are, and Degree of Agreement (DOA) [189], which quantifies the level of concordance between the ranking of items produced by a recommendation system and the ranking preferred by the user. Moreover, business metrics such as Click-Through Rate (CTR) [190] are critical to assessing the system’s impact on user engagement and the company’s bottom line, and Conversion Rate (CVR) [190] measures how efficient an algorithm is at providing recommendations that lead to user purchases. There also are some metrics for measuring other functionality. For example, Gini Index [191] evaluates the fairness of recommendation distribution, with lower values indicating more equitable distribution across items.
VIII Conclusion
This study comprehensively examines the integration of FRSs with FMs, a direction that has gained attention for its ability to protect user privacy. The article begins by summarizing common approaches of FRSs and FMs, then delves into the challenges faced during integration. To address these challenges, the paper proposes various strategies including using transfer learning and multitask learning techniques to adapt to data diversity, employing privacy-preserving methods like differential privacy and secure multi-party computation, and reducing communication overhead through model compression and efficient communication protocols.
This work also discusses the future research directions directions indicating that FRSs can provide more accurate personalized recommendations while better protecting user privacy. Additionally, the paper showcases applications of FRS in various fields, demonstrating their potential and value in the real world. Through this study, we aim to provide theoretical guidance for integrating FRS with FM, directing future research and technological advancements to collectively advance this field.