### v3

# Navigating the Future of Federated Recommendation Systems with Foundation Models

Abstract
In recent years, the integration of federated learning (FL) and recommendation systems (RS), known as Federated Recommendation Systems (FRS), has attracted attention for preserving user privacy by keeping private data on client devices. However, FRSs face inherent limitations such as data heterogeneity and scarcity, due to the privacy requirements of FL and the typical data sparsity issues of RSs. Models like ChatGPT are empowered by the concept of transfer learning and self-supervised learning, so they can be easily applied to the downstream tasks after fine-tuning or prompting. These models, so-called Foundation Models (FM), focus on understanding the human’s intent and perform following their designed roles in the specific tasks, which are widely recognized for producing high-quality content in the image and language domains. Thus, the achievements of FMs inspire the design of FRS and suggest a promising research direction: integrating foundation models to address the above limitations. In this study, we conduct a comprehensive review of FRSs with FMs. Specifically, we: 1) summarise the common approaches of current FRSs and FMs; 2) review the challenges posed by FRSs and FMs; 3) discuss potential future research directions; and 4) introduce some common benchmarks and evaluation metrics in the FRS field. We hope that this position paper provides the necessary background and guidance to explore this interesting and emerging topic.
Index Terms:
Federated Recommendation System, Foundation Model, Survey, Privacy Preserving, Security
I Introduction
In the digital age, the exponential growth of information has created a need for systems that navigate, filter, and personalize data for individual users. Recommendation Systems (RS) have become crucial tools for filtering online information and helping users discover products, content, and services that align with their preferences [1]. However, these systems traditionally rely heavily on centralized data collection and processing, posing significant privacy risks and operational bottlenecks. User privacy has become increasingly important, especially with stringent data protection regulations such as the General Data Protection Regulation (GDPR) [2], emphasizing storing user data locally rather than uploading it to servers.
To address these privacy concerns, Google introduced Federated Learning (FL) [3], designed to train models across distributed devices while keeping data localized. This paradigm shift in data processing leverages the computational capabilities of individual devices for local data analysis. Specifically, FL alternates between training models locally on each client and aggregating parameters globally on a central server. The integration of FL with RS is essential for safeguarding user privacy in recommendation services, giving rise to the burgeoning field of Federated Recommendation Systems (FRSs) [4].
FRSs have shown promising results in various areas, such as service provision [5, 6], daily scheduling [7], and driving planning [8, 9], impacting multiple facets of daily life. However, in most cases, each client typically consists of a single user’s device, containing only the data of one user’s accessed items, which is extremely small compared to the total item set, creating a serious data scarcity problem. In addition, different users have different behaviors and preferences, leading to significant data heterogeneity. The presence of both issues can result in sub-optimal models and reduced effectiveness. Addressing these challenges is crucial for the success of FRSs, particularly as they become more widely adopted in privacy-sensitive domains. Innovative strategies to enhance data sharing and model personalization while maintaining privacy are needed to overcome data scarcity and heterogeneity.
Figure 1: The foundation model can integrate information contained in data from various modalities during pre-training. The model can then be adapted for a variety of downstream tasks through adapters such as prompting or fine-tuning.
More recently, a novel paradigm for constructing artificial intelligence (AI) systems, known as Foundation Models (FMs) [10] (or Pretrained Large Models [11]), has garnered considerable interest following the remarkable success of models like ChatGPT [12] and various diffusion models [13, 14, 15]. As illustrated in Fig. 1, this paradigm involves using self-supervised learning objectives to pre-train models on large amounts of unlabeled data. This process, referred to as Pre-Training, enables the models to learn general-purpose representations. Once pre-trained, FMs can be adapted to specific tasks through techniques such as Fine-Tuning [16] or Prompting [17], turning the general model into a task-specific one. This adaptability allows FMs to achieve state-of-the-art performance across a wide range of applications. Examples of FMs include language models like BERT [18] and vision models like ViT [19], which have significantly advanced natural language understanding and computer vision, respectively. Multi-modal models like CLIP [20] and DALL·E [21] extend these capabilities to handle both text and images, enabling tasks such as image-text retrieval and text-to-image generation.
The widespread adoption and success of FMs demonstrate their profound impact on AI research and applications. FMs are typically trained on substantial quantities of data from diverse sources, enabling them to incorporate extensive knowledge when tackling specific downstream tasks [10]. This characteristic offers several significant benefits for FRSs [22]: Firstly, FMs can alleviate the data sparsity issue by providing rich pre-trained representations, thus enhancing recommendation accuracy even with limited user interactions. Secondly, they effectively mitigate the cold-start problem for new users and items by leveraging learned general patterns. Thirdly, FMs facilitate efficient transfer learning, allowing models to adapt quickly to new domains or tasks with minimal additional training. Moreover, the use of FMs can help balance privacy and performance; their pre-trained capabilities reduce the need for extensive data sharing and multiple rounds of communication in federated settings, thereby lowering the computational and communication costs associated with federated training. Given these considerable advantages, integrating FMs into FRSs is not just a viable approach but a necessary progression to overcome existing limitations and drive innovation in the field.
Although integrating FMs into FRSs holds immense promise, it must be acknowledged that this field is still in its nascent stages, with an insufficient understanding of the challenges, viable methods, and directions for development. This paper aims to bridge this knowledge gap through an in-depth analysis of the integration of FMs and FRSs. We provide a comprehensive examination of the motivations and challenges associated with uniting these two paradigms, drawing on insights from recent studies, which highlight the benefits of introducing FMs in federated recommendation scenarios. Additionally, we outline future development trends and potential applications. By elucidating the intersection of FRSs and FMs, this paper aims to promote further exploration and innovation within this emerging field, thereby facilitating its rapid advancement. The paper also emphasizes the critical role of pre-trained representations in addressing challenges like data sparsity and heterogeneity, which are inherent in FRSs.
Paper Organization. The reminder is structured to provide a thorough study of the current research landscape, technical challenges, application cases, and future directions of FRSs. The organization is as follows: Sec. II outlines the existing surveys within the field of FRSs, and distinctly highlights our novel contributions, setting our paper apart from the current literature. Sec. III introduce foundational knowledge in the field of FRSs and FM, including definitions, basic architectures, and current scenario taxonomy. Sec. IV introduces the potential applications and functions of FM according to the three classical phases of FRSs. Sec. VII summarizes the datasets and evaluation methods commonly used in FRSs with FM, explaining how to quantify the recommendation performances. Sec. V delves into the main technical challenges faced by FRSs with FM, including data heterogeneity, model aggregation difficulties, and privacy concerns, as well as current methods and their limitations in addressing these challenges. Sec. VI explores future research directions in FRSs, including unresolved issues and emerging research areas, providing guidance and inspiration for researchers in the field. Sec. VIII concludes our paper with a summary of key findings and perspectives, offering insights into the trends and potential future of FRSs.
II Related Surveys and Contribution
In the field of FRSs, several surveys have been conducted to synthesize methodologies, privacy preservation techniques, and inherent challenges, each emphasizing different aspects. Yang et al. [23] discuss the practical implementation and evaluation of FRSs, focusing on system architecture and algorithmic efficiency. Alamgir et al. [24] provide a comprehensive overview of FRSs, highlighting techniques, prevailing challenges, and future directions. Javeed et al. [25] focus on security and privacy concerns in personalized RSs, proposing targeted solutions to these challenges. Finally, Sun et al. [26] conduct a survey comparing existing federated recommendation research, highlighting the strengths and limitations of various approaches. Collectively, these surveys emphasize the critical importance of privacy protection and address challenges posed by data heterogeneity and model aggregation in FRSs. This work enhances our understanding of FRSs and provides a foundation for future research to tackle challenges.
While several studies have explored the combination of FMs with FL [27, 28, 29, 30, 31, 32, 33] or RSs [34, 35], to the best of our knowledge, no existing work has systematically analyzed or discussed the integration of FMs into FRSs. Our paper aims to fill this gap by providing a comprehensive examination of this intersection.
Contributions. Our main contributions are as follows:
  • We establish a comprehensive framework for applying FMs to FRSs, elaborating on the underlying principles and methodologies for effective integration.
  • By thoroughly analyzing existing technologies, we inspire innovation in leveraging the pre-training capabilities of FMs within FRSs, highlighting potential advancements.
  • We delve into the challenges and practical issues of integrating FMs into FRSs, including privacy protection, data heterogeneity, communication efficiency, and model generalization, providing insights and potential solutions.
  • We identify current research gaps and outline future directions, aiming to guide subsequent academic research and technological development in this emerging area.
  • Through application studies, we illustrate how the integration of FMs with FRSs can be applied in real-world scenarios, bridging the gap between theory and practice.
By elucidating the integration of FMs and FRSs, our paper seeks to promote further exploration and innovation, facilitating rapid advancement and broader adoption in the field.
Figure 2: The taxonomy of FRS and FM frameworks categorized by their respective core criteria in Sec. III.
III Background and Preliminary
To provide a comprehensive understanding of the integration of FRSs and FMs, this section outlines their core principles, taxonomies, and foundational structures, as illustrated in Fig. 2, establishing a basis for analyzing their synergies.
III-A Federated Recommendation System
Figure 3: A framework for FRSs, illustrating client-side local training with private data, server-side global aggregation, and update communication. Local models integrate user and item embeddings with prediction networks, while the server aggregates the sent local updates to enhance the global model, ensuring privacy and personalization.
Definition 1 (Federated Recommendation System)
A Federated Recommendation System is a technology that uses distributed algorithms for personalized information filtering. It aims to improve the accuracy of information filtering and the effectiveness of personalized services, while maintaining user privacy.
FRSs enable collaborative learning across multiple clients without direct exchange of raw user data. Fig. 3 illustrates that FRS typically involves three key processes: 1) client model update, 2) communication, and 3) global aggregation. Users train models locally using their own data and share updated parameters (e.g., gradients) with a server. The server then performs global aggregation, integrating information from all participants and redistributing updated parameters. This communication process preserves privacy and security.
(a) Horizontal FRS
(b) Vertical FRS
(c) Transfer FRS
Figure 4: Typical data distribution paradigms in FRS: (a) Horizontal FRS, characterized by similar feature spaces across users but distinct data samples; (b) Vertical FRS, involving overlapping user sets with disparate feature spaces; and (c) Transfer FRS, leveraging transfer learning to enable knowledge sharing across domains with limited or complementary data distributions.
As illustrated in Fig. 2, FRSs can be categorized based on data distribution and communication architecture [23, 26].
III-A1 Data Distribution Paradigms in FRS
There are three categories based the data distribution, each addresses specific data collaboration and learning scenarios:
  • Horizontal FRS deals with similar feature spaces but different sample spaces, as shown in Fig. 4a. It aggregates model updates from multiple clients to train recommendation models while preserving privacy [36, 37, 38].
  • Vertical FRS is applied when feature spaces differ but sample spaces are similar, as shown in Fig. 4b. It securely integrates different data types (e.g., a bank’s credit history) to enrich user information [39, 40, 41, 42].
  • Transfer FRS utilizes transfer learning [43] to transfer knowledge from one domain to another, benefiting scenarios with limited data for independent model training [22], shown in Fig. 4c.
Each type of FRS based on data distribution is designed to address specific collaboration scenarios, ensuring privacy and efficient learning. Horizontal FRS is suitable for situations where similar features are available across clients, Vertical FRS integrates different types of features for the same users, and Transfer FRS allows knowledge transfer across domains, making it particularly useful when data is scarce.
(a) Centralized FRS
(b) Semi-decentralized FRS
(c) Decentralized FRS
Figure 5: Architectural paradigms of FRS: (a) Centralized FRS, relying on a central server for model aggregation; (b) Semi-decentralized FRS, incorporating intermediate nodes to balance scalability and efficiency; and (c) Decentralized FRS, enabling peer-to-peer communication to enhance privacy and robustness.
III-A2 Communication Architectures in FRS
As shown in Fig. 5, Based on communication architecture, FRS can be categorized into centralized, semi-decentralized and decentralized based on their communication architecture. Each type addresses privacy and scalability in different ways:
  • Centralized FRS: A central server orchestrates the learning process in Fig. 5a, simplifying model aggregation to overcome data silo issues [44] while relying on a central authority. This approach is straightforward but depends heavily on the central server, creating a single point of failure and potential privacy risks [45, 36, 37, 38, 22].
  • Semi-decentralized FRS: As shown in Fig. 5b, It introduces an intermediate layer (e.g., edge servers) to reduce communication overhead and enhance scalability. This architecture balances the advantages of centralized and decentralized systems, improving system efficiency while still managing privacy concerns [46].
  • Decentralized FRS: It employs peer-to-peer communication [47], which maximizes privacy but necessitates complex coordination. As illustrated in Fig. 5c, each client communicates directly with others without relying on a central server. This enhances privacy and reduces dependence on a single entity; however, it comes at the expense of increased communication complexity and the requirement for robust consensus mechanisms [48, 49].
Each of these architectures offers a trade-off between privacy, communication efficiency and the degree of decentralization. Centralized systems simplify model aggregation, but rely on a central authority. Semi-decentralized systems aim to balance efficiency and privacy with an intermediate layer, while decentralized systems offer the highest level of privacy at the cost of more complex model coordination.
III-B Foundation Model
The rapid increase in the performance of computer hardwares, e.g., GPUs, the increasing maturity of transformer architectures, and the public availability of large amounts of training data have been three key factors in the emergence of FMs [10]. According to the work of Stanford HAI [10], we have the following definition of FMs:
Definition 2 (Foundation Model)
A foundation model is defined as any model trained on extensive data (typically using large-scale self-supervised learning) that is capable of adapting to a wide range of downstream tasks, for instance, through fine-tuning or prompting.
FMs are pre-trained on extensive data, enabling adaptation to a wide range of downstream tasks through fine-tuning or prompting [50, 51]. These models, e.g. GPT-3 [52], are distinguished by two key characteristics [10]: 1) Emergence, refers to the model’s ability to develop new capabilities from training data without explicit programming; and 2) Homogenization, highlights the unified approach across diverse machine learning tasks, making FMs adaptable across multiple domains.
Although FMs are based on self-supervised learning and transfer learning, their scale brings new emergent abilities, and their effectiveness in many tasks has motivated homogenization, e.g., many FMs have evolved based on transformer architectures [18, 53, 19, 52]. While homogenization provides strong leverage, it also requires caution, as defects in foundation models can be inherited by all downstream adapted models. Despite the broad deployment prospects of foundation models, due to their emergent properties, our understanding of how they work, when they fail, and what they can actually do is still quite limited [10], but there has been quite a bit of work [54, 55, 56] that attempts to explain this phenomenon. Moreover, OpenAI proposes a Scaling Law for LLM [57], suggesting that the performance of a model is governed by a power-law relationship involving computational power, model parameters, and data size. This law also applies to downstream tasks [58]. Therefore, understanding and leveraging this power-law relationship when using FM can aid in more efficiently allocating resources and optimizing model performance.
III-B1 Taxonomy of FM
As shown in Fig. 2, similar to FRS, FMs can also be classified based on data type or functionality. Specifically, based on the type of data used during training, FM is mainly divided into followings [10]:
  • Language FM: Trained on textual data to understand and generate language [18, 53], excelling in natural language processing tasks such as machine translation [59], search [60, 61], and sentiment analysis [62]. Typical language models include BERT [18], GPT-3 [52], and LLama [63], which typically use deep transformer architectures that are able to capture complex linguistic regularities and demonstrate excellent performance on multilingual tasks.
  • Vision FM: Focused on understanding image data, these models, such as DINOv2 [64] and SAM [65], are used for object recognition [19, 66, 67], visual reasoning [68, 69] and image segmentation [65, 70, 71].
  • Multi-modal FM: These models, e.g. CLIP [20] and DALL·E [21], integrate multiple types of data, such as text and images, allowing them to handle cross-modal tasks like image captioning and visual quizzing [72, 73, 74] by integrating information from different modalities.
Functionally, FMs can be categorized into [10, 75, 35]:
  • Discriminative FM: Primarily used for tasks like classification [76, 77] and regression [78], focusing on distinguishing between different classes of input data.
  • Generative FM: Such as GPT-3 and DALL·E, are designed to generate new data instances by learning the distribution of the training data [79], supporting applications like text [80] and image generation [14, 15].
Each type of FM offers unique advantages in its respective area. Language, vision, and multi-modal models provide specialized capabilities for processing different data types, while discriminative and generative models offer distinct approaches for solving problems based on either classification or data generation. By leveraging these specialized models, FMs can address a wide range of challenges across diverse domains.
III-B2 Adaptation
Fig. 1 illustrates a commonly used technique for pre-trained models, known as the Adaptation. Adaptation is a commonly used technique for pre-trained models, where new lightweight layers are added while maintaining the original parameters of the pre-trained model. This enables fine-tuning and extension for specific tasks [81, 38], making it suitable for multi-task learning and tasks performed in resource-limited environments, such as few-shot learning [82]. Adaptation techniques have found wide applications in natural language processing and computer vision tasks [83, 84, 85].
Adaptations can be categorized into three main types [82]:
  • Prompt-based Fine-tuning Adaptation Methods: These methods [86, 87] use prompt learning to adapt models for specific tasks by fine-tuning a portion of the fixed text input as a learnable vector. This allows the model to adjust quickly for new tasks with minimal changes.
  • Adapter-based Fine-tuning Adaptation Methods: Small, trainable modules, called Adapters, are added to a pre-trained model. These adapters [88, 85, 83] are fine-tuned instead of the entire model, making task adaptation efficient with minimal parameter updates.
  • External Knowledge-based Adaptation Methods: External knowledge, such as information from knowledge graphs or additional model queries, is integrated to improve the model’s performance on domain-specific tasks, especially when data is limited [89, 90].
In summary, adaptations play a crucial role in enhancing the versatility of foundation models. By enabling efficient fine-tuning and leveraging external knowledge, these methods help foundation models effectively tackle a wide range of tasks, even in resource-constrained settings or when training data is scarce. This adaptability is key to making foundation models more practical and powerful for real-world applications.
IV Federated Recommendation Systems with Foundation Models
Figure 6: Integration framework of FRSs with FMs, illustrating the key challenges in three stages: Client Model Update, Communication, and Global Aggregation.
As illustrated in Fig. 6, a typical FRS comprises three key stages: client model update, communication, and global aggregation [24]. The integration of FMs aligns with these stages. This section provides a comprehensive overview by summarizing representative works at each stage and examining additional dimensions of integrating FMs into FRS.
IV-A Client Model Update
In FRS, clients have the following characteristics [23, 24]:
  • Data Sparsity: As mentioned earlier, a major advantage of FRS is privacy. The data required for model training in FRS, such as user information and interaction history, is privacy sensitive and therefore must remain on the user’s device, creating data silos. This approach not only complies with data privacy regulations (such as GDPR), but also increases user trust. However, this means that each user has a relatively small amount of data compared to the entire dataset, and each user only accesses a portion of the item set, leading to data sparsity.
  • Non-IID Data: Each client independently trains the model on its local data. This allows the model to better adapt to each user’s unique preferences and behaviors, improving personalization and accuracy of recommendations. The resulting data distribution often does not satisfy the independent and identically distributed (IID) assumption, leading to the problem of data heterogeneity.
  • Device Limitations: In addition, user devices are typically consumer products such as mobile phones and personal computers, which are characterized by unstable communication and limited computational resources. This requires keeping the computational load of the models deployed on the client and the amount of information exchanged with the server as low as possible.
FMs are typically pre-trained on large, diverse datasets, providing them with rich knowledge and strong representational capabilities. This pre-training endows them with prior knowledge that can be quickly adapted to specific client data through fine-tuning [10]. Therefore, FMs learn general feature representations, offering adaptability to different data distributions. Applying FMs during the client update phase in FRS can significantly enhance the performance of local models. Specifically, the advantages are outlined as follows:
  • Addressing Data Sparsity: FMs can apply the knowledge learned from large corpora to local data through transfer learning [91]. For instance, a pre-trained FM can be effectively fine-tuned on a small amount of local data at the client to adapt to specific user behaviors, thereby achieving good performance in downstream tasks such as providing more accurate recommendations [35, 92]. Furthermore, by fine-tuning the base model locally, sensitive data does not need to leave the device, protecting user privacy [93]. Thus, even with limited data, clients can achieve better model performance, alleviating the challenges brought by data sparsity.
  • Handling Non-IID Data: FMs’ strong representational capabilities enable them to capture complex user preferences and behaviors. For example, FMs possess powerful semantic understanding abilities, allowing them to better interpret user search queries [60, 61], comments [34], and other textual data [94], thereby providing recommendations that align more closely with user needs. Moreover, FMs can learn complex patterns from non-IID data [11, 95, 96], enhancing the model’s adaptability to different user preferences and ensuring that recommendation results stay consistent with the user’s current interests and needs. Applying FMs during the client update phase can achieve a higher level of personalized recommendation, effectively addressing data heterogeneity.
  • Adapting to Device Limitations: Although fine-tuning and updating FMs typically require significant computational resources, optimizing the model architecture and employing lightweight fine-tuning techniques can reduce the computational burden on devices [97]. The knowledge transfer capability of FMs allows for efficient local fine-tuning under limited computational resources, thereby achieving good recommendation performance even in resource-constrained environments [98, 97]. Additionally, model compression techniques can be used to reduce the model size, thereby lowering computational and storage requirements to suit the limitations of user devices [99].
These characteristics make FMs highly suitable for addressing the diverse data requirements of clients in FRS, thereby improving the generalization performance of local models.
IV-B Communication
Communication between clients and the server is a critical phase in FRSs, involving the transmission of model update information from each client to the central server. Ensuring efficiency and security during this process is essential for maintaining system performance and protecting user privacy. However, this phase faces significant challenges, including the high communication costs associated with transmitting large volumes of data and the inherent security risks of handling sensitive information. Below, we firstly outline these challenges as follows:
  • Communication Costs: The transmission of large-scale model parameters, gradients, or other statistical data can result in excessive communication overhead. This is especially problematic in systems with many clients or high-dimensional models, where bandwidth and latency constraints become bottlenecks.
  • Transmission Security: Sensitive data encapsulated in model updates may be vulnerable to interception during transmission. Such breaches could lead to privacy violations or exposure of user data, undermining the privacy-preserving goals of FRS. Although encryption techniques can mitigate these risks, they often increase the computational and communication burden.
Addressing these challenges requires innovative approaches that enhance both efficiency and security. FMs, with their advanced representation and compression capabilities, offer promising solutions to overcome these limitations:
  • Reducing Communication Costs: FMs can perform semantic compression [100, 101], extracting critical features while eliminating redundant information, thus significantly reducing the data transmitted between clients and the server. Furthermore, optimized communication protocols tailored to FM representations [102] can further improve data exchange efficiency.
  • Enhancing Transmission Security: Transmitting sensitive data in FRSs poses significant privacy risks. FMs enhance transmission security by transforming raw model updates into abstract representations that retain essential information while removing sensitive details, reducing the risk of privacy leakage [33]. Additionally, FMs enable semantic compression [103], which not only reduces communication overhead but also minimizes the risk of adversarial attacks. When combined with techniques such as secure aggregation [104], FMs provide an efficient and scalable solution that strengthens both the security and performance of data transmission in FRS.
IV-C Global Aggregation
In FRS, the global aggregation phase of the server is responsible for combining model updates from all clients into a single global model. This phase is of critical importance in ensuring the overall performance and effectiveness of personalized recommendations. Server global aggregation is characterized by the following key features [105]:
  • Diverse Client Updates: The server must handle updates from a large number of clients, each containing different data and user behaviors. To ensure the overall quality of the model and the accuracy of recommendations, the server must effectively integrate this diverse data.
  • Noisy Data: The global aggregation process must be sufficiently robust to handle potentially noisy data and anomalous updates from clients, which can adversely affect the global model’s performance.
  • Scalability: The scalability of the server’s global aggregation is critical, especially when dealing with a large number of clients and massive amounts of data. The system must efficiently process and aggregate updates to maintain performance.
To address these challenges, several existing strategies have been developed. For example, in each iteration, given resource constraints and privacy concerns, the central server may only communicate with a subset of clients. Studies [106, 107] indicate that clients should not participate in successive training rounds, as attackers could potentially infer information about the client from transmitted gradients or model parameters. The client selection strategy determines which clients will participate in the current training round based on factors such as computational power, network stability, data diversity, and data quality [108]. For example, client selection can be done by random sampling or based on statistical characteristics of the data to ensure that the model learns from diverse data sources, thereby increasing its generalization ability. However, these existing strategies have limitations, such as suboptimal aggregation efficiency and challenges in fully exploiting the contextual information present in client updates.
FMs can play a significant role in enhancing the global aggregation process. Traditional aggregation methods, such as naive weighted averaging proposed in FedAvg [3], may not fully exploit the contextual information present in the data. FMs, however, are capable of intelligently aggregating updates by understanding the semantics and contextual relationships of each participant’s updates [70].
  • Context-Aware Aggregation: FMs utilize their powerful representation capabilities to analyze the structural and statistical characteristics of client updates, including update frequency, scale, and trends. This enables the server to more effectively assess the significance of each update without relying on potentially sensitive user-specific information, thereby aligning the aggregation process more closely with system-wide patterns and requirements [29]. By understanding the latent relationships between updates, FMs assign appropriate weights to individual contributions, improving the fidelity of the global model.
  • Dynamic Weighting: FMs enable dynamic adjustment of update weights based on their relevance to the global model [97]. For instance, updates contributing to under-represented data distributions may receive higher weights, ensuring that the global model remains inclusive and robust to diverse data sources [109].
  • Knowledge-Based Aggregation: Leveraging the extensive knowledge base of FMs [110], knowledge-based aggregation strategies have the ability to capitalize on FMs’ knowledge and reasoning capabilities to enhance the aggregation process. For instance, if certain clients lack sufficient data, FMs can use pertinent information from their knowledge base to infer relationships and patterns, thereby supplementing and improving the aggregation process to have better performance [111].
  • Anomaly Detection and Handling: FMs also excel in anomaly detection [112, 113]. Client updates may be influenced by various factors, leading to noisy or abnormal data [26]. If these abnormal updates are directly used in global aggregation, they may affect the overall model quality. FMs can analyze client updates to identify and handle anomalies, ensuring the robustness of global aggregation. For example, if a client’s data update shows a notable divergence from the majority, it may be regarded as noise or an anomaly, and FMs can choose to disregard or assign a reduced weight to it, thereby maintaining the stability of the global model [34].
Overall, FMs effectively tackle the key challenges in the global aggregation phase by integrating client updates with context-awareness, dynamically weighting to handle noisy data, leveraging pre-trained knowledge for scalability, and performing robust anomaly detection. These approaches significantly enhance the global model’s quality, robustness, and adaptability, positioning FMs as a promising tool for advancing FRSs. Continued research in this area could further unlock their potential, fostering progress in personalized, scalable, and privacy-preserving recommendations.
V Key Challenges and Strategic Solutions
Figure 7: Challenges in integrating FMs into FRSs, divided into data and model challenges, highlighting areas requiring strategic solutions for enhanced personalization and privacy.
Integrating FMs into FRSs has introduced new opportunities for enhancing personalized recommendations while maintaining user privacy. However, as illustrated in Fig. 7, this integration presents significant challenges across multiple dimensions. Understanding these challenges comprehensively and developing effective strategies to address them is crucial for advancing FRSs. This section explores these challenges in depth and presents potential strategies to overcome them.
V-A Data Challenges and Mitigation Strategies
V-A1 Privacy Concerns
Privacy aims to prevent unauthorized access or misuse of data, especially during model training and data handling, to protect user identities and personal details [114]. In a federated setting, safeguarding user privacy is of utmost importance. Privacy concerns focus on preventing unintended exposure or misuse of personal data, ensuring that sensitive information remains confidential. While training and recommendations must occur without revealing personal data, FMs, such as GPT-3, may memorize and reproduce training data, potentially leaking sensitive information [115]. Additionally, if generated data closely resembles original data, user privacy risks may arise [114].
To mitigate these privacy concerns, differential privacy [116] can be employed by adding noise during model training to reduce data leakage risks while preserving model performance. Homomorphic encryption [117] allows training on encrypted data, effectively preventing data theft during transmission and processing. Moreover, machine unlearning techniques [118] can be utilized to remove specific user data from the FMs, ensuring compliance with privacy regulations like GDPR. An emerging approach, known as privacy rewrite, transforms sensitive data into anonymized formats before transmission. Frameworks like HaS [115] exemplify this by anonymizing private entities locally and reconstructing them after processing. This method reduces privacy risks, preserves data utility, and minimizes computational overhead.
V-A2 Security Threats
Security threats involve protecting data from malicious attacks or breaches [114]. While privacy controls who can access and use data, security emphasizes maintaining data integrity and defending against threats such as cyberattacks [119]. The integration of FMs into FRSs also necessitates ensuring data security, as participants may be subject to attacks like Member Inference Attacks [119], Data Reconstruction Attacks [120], and Poisoning Attacks [121], which can compromise model integrity or data quality.
To ensure data integrity and confidentiality during transmission, secure multi-party computation [122] can distribute computational tasks among multiple participants, preventing any single party from accessing complete data. Additionally, blockchain [123, 124] can record data access and operations, enhancing data integrity, traceability, and transparency.
V-A3 Data Scarcity
Despite FMs’ potential to alleviate data sparsity, FRSs often face data scarcity, especially in scenarios where participants have limited interaction data under federated Settings [24]. Data sparsity can significantly impact model performance, particularly when user interactions are insufficient to build effective models.
Data augmentation techniques [125, 126] can be used to generate diverse training samples, expanding the dataset size. Knowledge transfer [98] from rich datasets in other domains can also alleviate data sparsity, improving overall model performance. When employing generative FMs [127, 128], quality control mechanisms are essential to ensure that synthetic data meets quality standards without introducing biases.
V-A4 Data Imbalance
In federated settings, significant differences in data size and distribution exist across clients, causing data imbalance [109, 129, 130]. This imbalance, often driven by long-tail distributions of item labels and user behaviors, affects model training effectiveness.
To address sample imbalance, resampling techniques like oversampling minority classes or undersampling majority classes [131] can be used. Weighted loss functions [132, 109] can also be used to assign higher importance to minority class samples during training, improving the capability to recognize underrepresented data without altering the data distribution.
V-B Model Challenges and Strategic Solutions
V-B1 Synchronization Overhead
In federated settings, the need for frequent synchronization between participants and the central server can result in high communication costs and system complexity, especially when dealing with large FMs with extensive parameters [10].
Gradient compression [133] can help reduce the data transmitted during synchronization, thereby lowering communication costs. Asynchronous update strategies [134] that allow participants to update models independently, followed by periodic global synchronization, also enhance system efficiency.
V-B2 Client Heterogeneity
Arising from variations in model types, sizes, and architectures, this issue creates challenges when deploying a unified FM in federated settings [135].
Adaptive training algorithms [136, 137] that adjust model parameters based on each client’s computational capabilities and data characteristics can address this issue. Knowledge transfer [22] between clients can further enhance model performance by mitigating disparities.
V-B3 Interpretability and Transparency
FMs are often perceived as black-box models, which makes it challenging to understand their internal workings and may lead to issues with trust and compliance [33].
Explainable AI [138], such as attention mechanisms [139] and feature importance analysis [140], offer insights into model decisions, enhancing transparency. Additionally, generative FMs [141] generate natural language explanations for recommendations, fostering user trust and acceptance.
Figure 8: Key future directions for integrating FMs into FRSs aim to address emerging challenges and opportunities, focusing on enhancing recommendation performance, safeguarding user privacy, and improving system adaptability.
VI Future Directions
The emerging field of applying FMs to FRSs presents both opportunities and challenges, shown in Fig. 8. By leveraging the powerful representation capabilities of FMs and the privacy-preserving advantages of FRSs, it is possible to significantly enhance recommendation performance while addressing some of the long-standing challenges faced by traditional methods. This section discusses several key future directions for integrating FMs into FRSs, focusing on specific problems, FMs’ advantages, the new challenges introduced, and potential research directions. Sec. VI aims to guide future research and applications towards delivering higher-quality recommendation services while protecting user privacy.
VI-A Data Augmentation
The challenge of data scarcity is a common problem in FRS due to data privacy constraints. Generative FMs, such as Large Language Models (LLMs) and Diffusion Models (DMs), can be utilized to alleviate this issue by generating synthetic user interaction data for training purposes, which can help mitigate data imbalance and scarcity [142, 143]. FM can generate virtual interaction records or additional item descriptions, thereby enhancing user profiles and improving recommendation accuracy and personalization [144].
However, the use of generated data introduces challenges related to data quality and potential biases, which may degrade the performance of recommendation systems. Rigorous validation and cleaning are necessary to ensure that the synthetic data is statistically consistent with real data, thereby reducing noise and biases [145]. Furthermore, integrating privacy-preserving techniques, such as differential privacy, can ensure that synthetic data generation does not reveal sensitive user information. Research should also focus on generating diverse data to cover a wide range of scenarios and user behaviors, thus reducing the risks of bias and overfitting [146].
VI-B Cold-Start Recommendation
The cold-start problem, caused by a lack of sufficient historical interaction data for new users or items, presents a significant challenge for FRSs [147]. FMs, pre-trained on large amounts of textual data, possesses rich semantic knowledge that can be leveraged for generating high-quality representations of users and items, enabling zero-shot and few-shot learning to address cold-start scenarios [148, 38].
Despite FMs’ capabilities, ensuring user privacy in a federated environment remains challenging. Additionally, multi-task learning techniques can be explored to improve model adaptability and generalization across different domains, allowing the model to provide effective recommendations for new users and items while maintaining privacy [149]. Future research should focus on enhancing privacy-preserving mechanisms and exploring efficient knowledge transfer techniques to support cold-start recommendations.
VI-C Multi-Modal Recommendation
The ability to process and integrate different modalities of data, such as text, images, audio, and video, is one of FM’s key strengths [10]. By incorporating multi-modal data, FM can construct more comprehensive user profiles [150], which in turn enhances the personalization of recommendation services [151, 152]. This richer profiling allows for more accurate user interest representation and better recommendation quality.
However, the heterogeneity of multi-modal data poses a challenge, as different data modalities have distinct structures and characteristics [153]. Developing unified representation methods to map different modalities into a common space is crucial for seamless model integration [150]. Privacy is also a concern, as multi-modal data often contains more sensitive information [154, 155]. Future research should focus on developing robust privacy-preserving techniques tailored for multi-modal data and improving computational efficiency in processing and integrating these diverse data types.
VI-D Real-Time Recommendations
Real-time recommendation is critical in FRS, as it allows for dynamically adapting recommendations based on real-time user behavior and contextual data [156]. FM’s contextual understanding capabilities enhance the accuracy and relevance of real-time recommendations by effectively processing user queries and item descriptions.
However, real-time recommendations require extremely low latency, which places a high demand on computational resources. FM’s high computational requirements may introduce latency issues, adversely affecting the user experience. Future research should focus on model compression and acceleration techniques, such as distillation [157] and pruning [158], to reduce computational complexity and improve response times. Additionally, methods like sliding windows can be used to manage the input context efficiently, optimizing the handling of long sequences of user behavior.
VI-E Enhanced Explainability
Explainability is an essential aspect of FRS, as providing transparent explanations for recommendations increases user trust and satisfaction. Language FMs, pre-trained on extensive text data, are capable of generating coherent and informative explanations that can help users understand the reasons behind recommendations [159, 160]. For example, models such as ChatGPT can generate natural language explanations for why specific items are recommended [161].
The challenge, however, lies in the computational cost associated with generating these detailed explanations, especially for large language models, which can lead to latency issues [10]. Moreover, biases present in pre-training data may lead to unfair or biased explanations [28]. Future research should focus on model compression to reduce computational costs and on debiasing methods to minimize biases in generated explanations. Incorporating user feedback may further help improve the quality and fairness of recommendation explanations.
VI-F Advanced Metrics
The evaluation of FRSs with FMs requires advanced metrics that extend beyond traditional metrics like rating prediction or item ranking [162]. Given FM’s ability to generate new items and provide explanations, traditional metrics may not fully capture its performance. New evaluation metrics that account for the generative and explanatory capabilities of FMs are needed to better assess recommendation quality in these emerging scenarios [35].
Future research should explore comprehensive evaluation criteria that consider aspects such as diversity, fairness, and user satisfaction to provide a holistic assessment of FRS performance with FM integration.
VII Resources
In this section, we introduce the commonly used datasets and evaluation metrics in FRS. The aim is to provide a comprehensive overview of the resources used in the field, helping readers understand the types of materials and benchmarks typically employed. This section presents various datasets, including explicit and implicit feedback types, along with metrics for evaluating recommendation performance. These datasets and metrics are crucial for assessing and improving the effectiveness of FRSs.
VII-A Datasets
Figure 9: Two types of Feedbacks.
TABLE I: Datasets with Explicit Feedback.
Context
Dataset #User #Item #Feedback Sparsity User Item Feedback Timestamp
Amazon Beauty 1,210,271 249,274 2,023,070 99.99%
Amazon Books 8,026,324 2,330,066 22,507,155 99.99%
Amazon CDs 1,578,597 486,360 3,749,004 99.99%
Amazon Cell Phone 2,261,045 319,678 3,447,249 99.99%
Amazon Clothing 3,117,268 1,136,004 5,748,920 99.99%
Amazon Games 826,767 50,210 1,324,753 99.99%
Amazon Garden 714,791 105,984 993,490 99.99%
Amazon Home 2,511,610 410,243 4,253,926 99.99% $\checkmark$ $\checkmark$
Amazon Kindle 1,406,890 430,530 3,205,467 99.99% $\checkmark$ $\checkmark$
Amazon Movies 2,088,620 200,941 4,607,047 99.99% $\checkmark$ $\checkmark$
Amazon Music 478,235 266,414 836,006 99.99% $\checkmark$ $\checkmark$
Amazon Sports 1,990,521 478,898 3,268,695 99.99% $\checkmark$ $\checkmark$
Amazon Toys 1,342,911 327,698 2,252,771 99.99% $\checkmark$ $\checkmark$
Amazon Reviews Amazon Video 426,922 23,965 583,933 99.99% $\checkmark$ $\checkmark$
ML-100K 943 1,682 100,000 93.70% $\checkmark$ $\checkmark$ $\checkmark$
ML-1M 6,040 3,952 10,000,209 95.81% $\checkmark$ $\checkmark$ $\checkmark$
ML-10M 69,878 10,681 10,000,054 98.69% $\checkmark$ $\checkmark$
MovieLens ML-20M 138,493 27,278 20,000,263 99.47% $\checkmark$ $\checkmark$
Yelp2018 1,326,101 174,567 5,261,669 99.99% $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
Yelp2020 1,968,703 209,393 8,021,122 99.99% $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
Yelp2021 2,189,457 160,585 8,635,403 99.99% $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
Yelp2022 1,987,929 150,346 6,990,280 99.99% $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
Yelp Yelp-full 5,556,436 539,254 28,908,240 99.99% $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
Anime 73,515 11,200 7,813,737 99.05%
Book Crossing 105,284 340,557 1,149,780 99.99% $\checkmark$ $\checkmark$
Douban 738,701 28 2,125,056 89.73% $\checkmark$ $\checkmark$
Epinions 116,260 41,269 188,478 99.99% $\checkmark$ $\checkmark$
Goodreads 876,145 2,360,650 228,648,342 99.99% $\checkmark$ $\checkmark$
Jester 73,421 101 4,136,360 44.22%
Netflix 480,189 17,770 100,480,507 98.82% $\checkmark$
Yahoo Music 1,948,882 98,211 11,557,943 99.99% $\checkmark$
TABLE II: Datasets with Implicit Feedback.
Context
Dataset #User #Item #Feedback Sparsity User Item Feedback Timestamp
MIND_large_train 711,223 27,047 83,507,374 99.57% $\checkmark$
MIND_large_dev 255,991 6,998 14,085,557 99.21% $\checkmark$
MIND_small_train 50,001 20,289 5,843,444 99.42% $\checkmark$
MIND MIND_small_dev 50,001 5,370 2,740,998 98.98% $\checkmark$
QK-video 5,022,750 3,753,436 142,321,193 99.99% $\checkmark$ $\checkmark$
QK-article 1,325,838 220,122 46,111,728 99.98% $\checkmark$ $\checkmark$
QB-video 34,240 130,647 1,701,171 99.96% $\checkmark$ $\checkmark$
Tenrec QB-article 24,516 7,355 348,736 99.81% $\checkmark$ $\checkmark$
Adressa 15,514 923 2,717,915 81.02% $\checkmark$ $\checkmark$
Foursquare 1,083 38,333 227,428 99.45% $\checkmark$ $\checkmark$
Gowalla 196,591 950,327 6,442,890 99.99%
Last.FM 1,892 17,632 92,834 99.72% $\checkmark$
Pinterest 55,187 9,911 1,445,622 99.74%
Steam 2,567,538 32,135 7,793,069 99.99% $\checkmark$ $\checkmark$ $\checkmark$
TaFeng 32,266 23,812 817,741 99.89% $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
Tmall 963,923 2,353,207 44,528,127 99.99% $\checkmark$ $\checkmark$
As shown in Fig. 9, the datasets used can be categorized into two types based on user feedback in FRS: explicit feedback and implicit feedback. Explicit feedback includes direct responses from users about their preferences. This typically includes ratings (such as a 1-5 scale), user comments, and like/dislike statements. Such data provides straightforward insights into user preferences, making it highly valuable for training recommendation models. Common datasets with explicit feedback include the following shown in Table I:
  • Amazon Reviews [^1] https://cseweb.ucsd.edu/ jmcauley/datasets.html#amazon_reviews: The Amazon Reviews dataset is a large-scale dataset that contains product information across various categories such as Books, CDs, and Music. It includes reviews (ratings, text, helpfulness votes) and product metadata (description, category information, price, brand, and image features). There are three updated versions of this dataset from the years 2014 [163, 164], 2018 [165], and 2023 [166].
  • MovieLens Datasets [^2] https://grouplens.org/datasets/movielens/: The MovieLens datasets [167], initially released in 1998, capture individuals’ stated movie preferences. These preferences are recorded as tuples, with each tuple showing a person’s rating (from 0 to 5 stars) for a movie at a specific time. Users enter these ratings through the MovieLens website, which provides personalized movie suggestions based on these ratings.
  • Yelp Datasets [^3] https://www.yelp.com/dataset: This dataset is a subset of Yelp’s business, review, and user data. It was initially developed for the Yelp Dataset Challenge, which allows students to study or analyze Yelp data and present their insights. In total, there are four versions of the Yelp datasets.
  • Anime [^4] https://www.kaggle.com/datasets/CooperUnion/anime-recommendations-database: It collects user preference data from the MyAnimeList website, and contains information from 73,516 users on 12,294 anime titles. Users can add anime to their completed list and rate them, and this dataset compiles these ratings.
  • Book Crossing [^5] https://grouplens.org/datasets/book-crossing/: The Book-Crossing dataset is a well-structured collection of data collected by Cai-Nicolas Ziegler in a 4-week crawl from the Book-Crossing community. This dataset primarily comprises user interactions that include book ratings, ranging from 0 to 10.
  • Douban [^6] https://www.kaggle.com/datasets/utmhikari/doubanmovieshortcomments: The Douban Movie dataset is a Chinese website where internet users can post their opinions and comments about films. It contains over 2 million short comments on 28 movies from the Douban Movie website.
  • Epinions [^7] https://cseweb.ucsd.edu/ jmcauley/datasets.html#social_data [168]: This dataset was collected from Epinions.com, a popular online consumer review site. It includes trust relationships between users and covers a period from January 2001 to November 2013.
  • Goodreads [^8] https://www.kaggle.com/datasets/jealousleopard/goodreadsbooks: This dataset includes reviews from the book review website Goodreads, along with various attributes describing the books. Importantly, the dataset captures different levels of user interaction, from adding books to a shelf, to rating them, to reading them.
  • Jester [^9] https://eigentaste.berkeley.edu/dataset/: It focuses exclusively on jokes. Users of the Jester online platform rate jokes and these ratings are then used to personalize joke recommendations for them.
  • Netflix [^10] https://www.kaggle.com/datasets/netflix-inc/netflix-prize-data: Netflix provided a training dataset consisting of 100,480,507 ratings from 480,189 users for 17,770 films. The user and the movie are identified by integer IDs, and the rating scores range from 1 to 5 stars, also as integers.
  • Yahoo Music [^11] https://webscope.sandbox.yahoo.com/catalog.php?datatype=r: It is known for its diversity and large scale. It contains a large collection of user ratings on different musical elements such as artists and genres.
Implicit feedback, on the other hand, is derived from user actions that indirectly indicate preferences, such as bookmarks, video/music play history, or click-throughs. Although implicit feedback does not directly express user likes or dislikes, it is rich and captures user behaviour more comprehensively.
  • MIND [^12] https://msnews.github.io/ [169]: The MIND dataset, sourced from the Microsoft News website, is a large-scale collection of approximately 160,000 English news articles and over 15 million user interaction records. It has been designed to advance research in news recommendation systems. It includes detailed textual content for each story and anonymized user interaction data to ensure privacy.
  • Tenrec [^13] https://github.com/yuangh-x/2022-NIPS-Tenrec [170]: The Tenrec dataset is a comprehensive benchmark dataset for RSs, featuring user interactions from two recommendation platforms across four dataset files: QK-video and QB-video for video actions, and QK-article and QB-article for article actions.
  • Adressa [^14] https://reclab.idi.ntnu.no/dataset/ [171]: It is a corpus of Norwegian news articles related to anonymous users. It is a collaborative project between the Norwegian University of Science and Technology and Adressavisen. The objective is to gain insight into the nature of news articles and their readers.
  • Foursquare [^15] https://sites.google.com/site/yangdingqi/home/foursquare-dataset [172]: This dataset comprises check-in data from New York City, collected over a period of approximately ten months (from 12 April 2012 to 16 February 2013). It encompasses 227,428 check-ins in New York City, with each check-in recorded with its timestamp, GPS coordinates, and detailed venue category.
  • Gowalla [^16] https://snap.stanford.edu/data/loc-gowalla.html [173]: Gowalla is a location-based social networking website where users can post their whereabouts by checking in. The dataset comprises data collected from the public API, which represents an undirected friendship network with 196,591 nodes and 950,327 connections. Additionally, it records 6,442,890 check-ins made by these users between February 2009 and October 2010.
  • Last.FM [^17] https://grouplens.org/datasets/hetrec-2011/ [174]: The Last.FM dataset represents a valuable resource that has been extensively utilized in the field of music information retrieval and RSs. It captures detailed information regarding music listening events from users. Each listening event is further enhanced with user demographics and specific descriptors that reflect their music tastes and consumption behaviours.
  • Pinterest [^18] https://github.com/edervishaj/pinterest-recsys-dataset [175]: The Pinterest dataset represents a valuable resource for a variety of research and analytical purposes. It encompasses a diverse range of data, including images, user features, interests, and user interactions.
  • Steam [^19] https://github.com/kang205/SASRec [176]: Steam is a collection of information about games published on the Steam platform. It includes details such as game names, release dates, genres, developers, publishers, and other relevant information.
  • TaFeng [^20] https://www.kaggle.com/datasets/chiranjivdas09/ta-feng-grocery-dataset: TaFeng is a comprehensive collection of supermarket shopping data, including detailed transaction records from the Ta Feng supermarket in Taiwan, covering a period from November 2000 to February 2001. The dataset comprises a variety of data points, including customer demographics, product categories, and detailed item descriptions along with quantities purchased.
  • Tmall [^21] https://tianchi.aliyun.com/dataset/53[177]: The Tmall dataset is a comprehensive collection from Tmall, comprising anonymized user shopping records over a six-month period up to and including the ”Double 11” event. It should be noted that the data is selectively sampled to address privacy concerns.
Both types of feedback play a critical role in the development of FRSs, providing diverse insights into user preferences that help improve the recommendation accuracy and relevance.
Figure 10: A potential unified framework for FRSs integrates server-side global aggregation and hyperparameter optimization with client-side data processing, model training, and evaluation.
VII-B Metrics
In the field of FRSs, the use of evaluation metrics is fundamental to assessing and refining the performance of our algorithms. Metrics serve as a quantitative lens through which we can observe how closely the system’s suggestions match users’ actual interests and preferences.
For predicting how well a system can estimate user preferences, there are several metrices measuring the prediction errors to judge the accuracy of the predictions [178], such as Mean Absolute Error, Mean Squared Error and Root Mean Squared Error. When it comes to classifying items, i.e., determining whether a user will like a product or not, we look at metrics such as Precision, Recall, Hit Ratio, F1 Score, Accuracy and AUC. These tell us how correctly the RS is classifying items, and the F1 score helps us balance the Precision and Recall. Then to measure the item ranking ability, which is about listing recommendations in the right order, Average Precision and Mean Average Precision are key to this, as they assess the quality of the order of recommendations. Metrics such as Mean Reciprocal Rank [179], Normalized Mutual Rank [180] and Normalised Discounted Cumulative Gain [181] also contribute by assessing how well the top recommended items are ranked.
On a broader scale, we consider recommendation-centric metrics such as Diversity, which ensures that a variety of items are suggested, and Coverage, which measures how many items from the catalogue are recommended. There are also some user-centric metrics include Novelty, which measures how new or surprising the recommendations are, and Degree of Agreement (DOA) [182], which quantifies the level of concordance between the ranking of items produced by a recommendation system and the ranking preferred by the user. Moreover, business metrics such as Click-Through Rate [183] are critical to assessing the system’s impact on user engagement and the company’s bottom line, and Conversion Rate [183] measures how efficient an algorithm is at providing recommendations that lead to user purchases. There also are some metrics for measuring other functionality. For example, Gini Index [184] evaluates the fairness of recommendation distribution, with lower values indicating more equitable distribution across items.
VII-C Open-source framework
Although there are already mature open-source frameworks for both RSs [185, 186, 187, 188] and FL [189, 190, 191], there remains a significant gap when it comes to FRSs. Currently, there is no mature, unified, and widely applicable open-source framework specifically designed for FRSs. Efforts such as FedMR [^22] https://github.com/mtics/FedMR [150] attempt to bridge this gap by integrating features from existing frameworks like MMRec [^23] https://github.com/enoche/MMRec [185] and RecBole [^24] https://github.com/RUCAIBox/RecBole [185], extending them into the realm of federated multimodal recommendation. However, FedMR is still in its infancy and does not yet provide the robustness or general applicability required of a comprehensive solution for FRSs.
As illustrated in Fig. 10, we make a preliminary attempt to propose a potential unified framework for FRSs in this paper, which aims to ideally encompass the following:
  • Integration of RSs and FL: Such a framework should seamlessly integrate the strengths of traditional recommendation systems with privacy-preserving capabilities from federated learning, addressing both data heterogeneity and privacy constraints effectively.
  • Modular and Extensible Architecture: The framework should feature a modular design that separates core components such as data preprocessing, model training, evaluation, and aggregation, facilitating improvements and the easy integration of new modules or functionalities.
  • Support for Multimodal Data: Leveraging the strengths of multimodal models, the framework should support the incorporation of various data modalities, enabling a more comprehensive understanding of user preferences and improving recommendation accuracy.
  • Scalability and Device Adaptability: The framework must be capable of scaling to accommodate a large number of clients and adapt to heterogeneous devices with varying computational capabilities, ensuring efficient performance on resource-constrained devices.
  • Privacy and Security: The framework should incorporate advanced privacy-preserving mechanisms, such as differential privacy and secure multi-party computation, to guarantee data confidentiality during both training and communication phases.
Despite the absence of a mature open-source FRS framework, the existing building blocks in RS and FL frameworks provide a strong foundation. By extending these capabilities and overcoming the unique challenges posed by FRSs, our proposed framework serves as a step towards a widely adopted, unified solution that fully supports FRSs.
VIII Conclusion
This paper comprehensively examines the integration of FRSs with FMs, a direction that has gained attention for its ability to protect user privacy. It begins by summarizing common approaches of FRSs and FMs, then delves into the challenges faced during integration. To address these challenges, we comprehensively various strategies to adapt to data diversity, employing privacy-preserving methods, and reducing communication overhead through model compression and efficient communication protocols.
This work also discusses the future research directions directions indicating that FRSs can provide more accurate personalized recommendations while better protecting user privacy. Additionally, we showcase applications of FRS in various fields, demonstrating their potential and value in the real world. Through this paper, we aim to provide a guidance for integrating FRS with FM, directing future research and technological advancements to collectively advance this field.